### v1

# Exploring the mysterious high-ionization source powering [NeĀ V] in high- analog SBS0335-052Ā E with JWST/MIRI

Abstract
Nearby blue compact dwarf galaxies (BCDs) share similar properties with objects from the Epoch of Reionization revealed by JWST, in terms of low stellar mass, low metallicity and high specific star-formation rate. Thus, they represent ideal local laboratories for detailed multi-wavelength studies to understand their properties and the mechanisms shaping them. We report the first JWST MIRI/MRS observations of the BCD SBSĀ 0335-052Ā E, analyzing MIR emission lines tracing different levels of ionization (e.g., [NeĀ II], [SĀ IV], [NeĀ III], [OĀ IV], [NeĀ V]) of the ionized gas. SBSĀ 0335-052Ā E MIR emission is characterized by a bright point source, located in one of the youngest and most embedded stellar clusters (Ā Myr, ), and underlying extended high-ionization emission (i.e., [OĀ IV], [NeĀ V]) from the surroundings of the older and less dusty stellar clusters (Ā Myr, ). From the comparison with state-of-the-art models, we can exclude shocks, X-ray binaries, and old stellar populations as the main sources of the high ionization. Interestingly, a 4-16% contribution of a Ā Mā intermediate massive black hole (IMBH) is needed to justify the strong [NeĀ V]/[NeĀ II]Ā and would be consistent with optical/UV line ratios from previous studies. However, even IMBH models cannot explain the strongest [OĀ IV]/[NeĀ III]. Also, star-forming models (regardless of including X-ray binaries) struggle to reproduce even the lower ionization line ratios (e.g., [SĀ IV]/[NeĀ II]) typically observed in BCDs. Overall, while current models suggest the need to account for an accreting IMBH in this high- analog, limitations still exist in predicting high-ionization emission lines (I.P.Ā Ā eV) when modeling these low-metallicity environments, thus other sources of ionization cannot be fully ruled out.
1 Introduction
In the JWST era, surveys such as the JWST Advanced Deep Survey (JADES), the Cosmic Evolution Early Release Science Survey (CEERS) and the Ultradeep NIRSpec and NIRCam Observations before the Epoch of Reionization (UNCOVER) have started to revolutionize our knowledge of the early Universe up to (e.g., Bagley etĀ al. 2023; Bunker etĀ al. 2023; Bezanson etĀ al. 2024; Carniani etĀ al. 2024a), exploring the epoch of reionization (EoR, ) as never done before. Several studies over the past decade have suggested that low-mass, metal-poor and highly star-forming galaxies can play a crucial role in the EoR (e.g., Wise etĀ al. 2014; Robertson etĀ al. 2015; Stark etĀ al. 2015), as now supported by recent JWST works probing the details of ultra-faint galaxies (e.g., Atek etĀ al. 2024; Simmonds etĀ al. 2024). JWST has also started to reveal an unprecedented number of active galactic nuclei (AGN) in the first billion years, comprising both intrinsically faint and reddened sources (e.g., Harikane etĀ al. 2023; Maiolino etĀ al. 2023a; LabbĆ© etĀ al. 2023; Kocevski etĀ al. 2024; Ćbler etĀ al. 2024; Juodāvzbalis etĀ al. 2024), including a new family of objects, the so-called Little Red Dots (e.g., Furtak etĀ al. 2023; Kokorev etĀ al. 2024; Matthee etĀ al. 2024). The role that these AGN can play in the EoR is currently under investigation (e.g., Dayal etĀ al. 2024), but it is still challenging to unambiguously identify AGN activity in EoR systems.
To detect the presence of faint AGN at , many works used the detection of broadened Balmer lines (FWHMĀ Ā km/s; e.g., Harikane etĀ al. 2023; Maiolino etĀ al. 2023a; Larson etĀ al. 2023; Kokorev etĀ al. 2023; Greene etĀ al. 2024; Greene & Ho 2004). However, the broad component under forbidden lines can simply be too faint to be detected (e.g., Carniani etĀ al. 2024b), leading to possible mis-classification of a pure star-forming system with a fast gas outflow as a broad line AGN. AGN classification has been also proposed via the presence of high-ionization rest-UV and blue optical lines (e.g., NĀ IV]Ā , [NĀ III]Ā , CĀ IVĀ 1548,51, [NeĀ IV]Ā 2423, HeĀ IIĀ 4686; Maiolino etĀ al. 2023b; Kokorev etĀ al. 2024; Greene etĀ al. 2024; Backhaus etĀ al. 2024). However, optical and UV diagnostics can be less effective in metal-poor systems (e.g., Feltre etĀ al. 2016; Ćbler etĀ al. 2023), because a hard stellar radiation field and AGN activity can behave similarly in low-metallicity dense environments (e.g., Leitherer etĀ al. 1999; Eldridge etĀ al. 2017; Telford etĀ al. 2023). Thus, metal-poor star-forming systems can show UV and optical high-ionization lines similar to those observed in high-z AGN candidates (e.g., the star-forming RXCJ2248-ID at , Topping etĀ al. 2024), making the two scenarios difficult to distinguish (e.g., Castellano etĀ al. 2024; Ćlvarez-MĆ”rquez etĀ al. 2024). Another proposed method to identify AGN is through photometric variability resulting from changes in AGN mass accretion rate (Hayes etĀ al. 2024), but larger samples of AGN at high- are needed to reduce the statistical and systematic uncertainties.
In this context, nearby galaxies can provide a precious opportunity for us to understand the characteristics of the interstellar medium (ISM) of these primeval objects and how to distinguish between star-formation and AGN activity within them, given the high signal-to-noise (S/N) of their spectra and the possibility to perform spatially resolved and multi-wavelength studies. In particular, nearby blue compact dwarf galaxies (BCDs) are ideal analogs to these high- systems, as they share properties in terms of stellar mass, intense star formation activity and metal poor content, as also suggested by recent JWST studies (e.g., Schaerer etĀ al. 2022; Brinchmann 2023). Many studies have explored these objects to understand how to constrain their ISM properties using their optical and UV emission (e.g., James etĀ al. 2009; Izotov etĀ al. 2012; Berg etĀ al. 2019; 2021; Senchyna etĀ al. 2021; 2023; Mingozzi etĀ al. 2022; 2024; Kumari etĀ al. 2024). In this context, the mid-infrared (MIR) wavelength range covers a large range of ionization potentials for the same species (e.g., 20-126Ā eV, [NeĀ II]Ā 21.56m, [NeĀ III]Ā 15.56m, [NeĀ V]Ā 14.32,24.3m, [NeĀ VI]Ā ) at lower excitation energies. While these lines are not accessible in the high- universe, they can provide an important constraint on the gas ionization mechanisms and, by also probing other phases (i.e., warm molecular gas with H2 rotational lines; dust with polycyclic aromatic hydrocarbons, PAHs), they can give a broad overview of the ISM conditions.
In this work, we focus on the galaxy SBSĀ 0335-052Ā E (; 1āĀ Ā pc), discovered by Izotov etĀ al. (1990). This object is a nearby BCD, extremely metal-poor (Ā Zā; 12+log(O/H)Ā ; Papaderos etĀ al. 2006) and characterized by a compact vigorous starburst (specific star formation rate, log(/Gyr-1)Ā ; RĆ©my-Ruyer etĀ al. 2015) and massive young clusters (Papaderos etĀ al. 1998; Reines etĀ al. 2008; Adamo etĀ al. 2010 and references therein). This galaxy is also well-known for a complex kinematics and a large scale HĀ outflow (Herenz etĀ al. 2023) and high-ionization optical and UV emission lines typically observed in the EoR (e.g., HeĀ IIĀ 4686, CĀ IVĀ , HeĀ IIĀ 1640; Herenz etĀ al. 2017; Kehrig etĀ al. 2018; Wofford etĀ al. 2021). In particular, its extended nebular HeĀ IIĀ 4686 emission requires energy beyond 54.4Ā eV (i.e., Ā Ryd, Ā Ć ), indicative of a very hard ionizing spectral energy distribution (SED), that only rotating metal-free or extremely metal-poor massive stars seem capable of reproducing (Kehrig etĀ al. 2018; see also Wofford etĀ al. 2021). Recent studies have also suggested that SBSĀ 0335-052Ā E may host an intermediate massive black hole (IMBH) of Ā Mā to explain its [NeĀ V]Ā 3426 emission (Hatano etĀ al. 2024, see also Thuan & Izotov 2005), requiring energies beyond 97Ā eV (i.e., Ā Ryd, Ā Ć ), as well as possible near IR variability (Hatano etĀ al. 2023). IMBHs (see Greene etĀ al. 2020 for a review) could represent the missing piece between stellar and supermassive BHs and would help explain the rapid accretion processes in place at high-. Thus, it is important to understand if galaxies such as SBSĀ 0335-052Ā E can host an accreting IMBH, and, if so, which are the best diagnostics to trace its activity and impact on the ISM.
To this end, we exploit JWST/MIRI Medium Resolution Spectroscopy (MRS) integral field spectroscopy (IFS) data to investigate the possible accreting IMBH or other non-stellar source ionizing the high- analog SBSĀ 0335-052Ā Eās ISM, using spatially resolved high ionization MIR lines. Indeed, SBS 0335-052Ā E intense SF activity makes this galaxy particularly bright in the MIR, which represents % of its total luminosity (Plante & Sauvage 2002). Spitzer IRS Ā m spectrum already revealed strong MIR emission lines (i.e., [SĀ IV]Ā 10.51, [NeĀ III]Ā 15.55, [OĀ IV]Ā 25.89), warm molecular gas features (i.e., H2Ā Ā S(3)) and the absence of PAHs (Houck etĀ al. 2004; Hao etĀ al. 2009). Nevertheless, MIRI/MRS offer the considerable advantage of spatially resolving the MIR emission (down to ā, Ā pc) and thus revealing the location of very high ionization extended emission (i.e., [NeĀ V]Ā 14.32m) for the first time in BCDs.
In Sec.Ā 2, we present the observations and data reduction. In Sec.Ā 3, we show the methods we adopted for our data analysis, including spectral fitting and literature photoionization and shock models that we used to interpret our results, reported in Sec.Ā 4. In Sec.Ā 5, we discuss our findings as well as possible limitations in current models, while in Sec.Ā 6, we conclude with a summary of our work. Throughout this work we assume a flat CDM cosmology (HĀ km/s/Mpc, ) and 12+log(O/H) (Asplund etĀ al. 2009).
2 Observations and Data Reduction
SBS 0335-052Ā E data were acquired as part of the Cycle 2 JWST program PID 4278 (PI: Mingozzi) on February 11, 2024. In particular, the galaxy was observed with JWST/MIRI MRS using the three MRS grating settings (SHORT, MEDIUM, and LONG; Ā ), and the FAST readout mode. Just one pointing was used since MIRI/MRS covers the majority of SBSĀ 0335-052Ā Eās active starburst activity with ChannelĀ 1 (3.2ā3.7ā) and fully samples it in ChannelĀ 4 (6.6ā7.7ā). Given the JWST pointing accuracy of ā, target acquisition was not required. A 4-point dither pattern was also applied to achieve optimal sampling throughout the MRS field of view (FOV) and to identify and remove detector artefacts. Each pointing was observed with 50 groups per integration for a total on-source exposure time of 1665Ā s. Dedicated background observations were performed with half the integrations (same group number but half the dithering) to accurately measure and correct for the thermal background. Background and science observations were connected via a fixed offset and placed in an uninterruptible sequence, and a PA orientation range was required to obtain simultaneous imaging observations using the F770W, F1130W, and F1800W filters, which will be presented in a forthcoming paper.
We retrieved MIRI/MRS raw products from Mikulski Archive for Space Telescopes (MAST) and we processed the data with v1.16.0 of the JWST calibration pipeline, following the jupyter notebook provided by the MIRI team. As extra steps to the standard version of the MIRI pipeline, we masked additional regions affected by cosmic showers at the end of StageĀ 1 (ShowerMasking Notebook) and we applied residual fringe corrections in StageĀ 2. We subtracted the background in StageĀ 3, using the master background created by the pipeline. In StageĀ 3 we created the final datacubes per band, to optimize the fringing correction, obtaining 12 datacubes centered to the same coordinates and keeping the native pixel scale size (i.e., 0.13ā, 0.17ā, 0.20ā and 0.35ā, for Channel 1, 2, 3 and 4, respectively). Nevertheless, we note that few spaxels located at the peak of the MIR emission are still affected by significant fringing, especially in ChannelsĀ 3 and 4. The FOV progressively increases from 3.2ā3.7ā in Channel 1, to 6.6ā7.7ā in Channel 4, as well as the MIRI/MRS PSF and spatial resolution, from in Channel 1 to in Channel 4.
In Fig.Ā 1 we show the position of MIRI/MRS FOV superimposed on the HST ACS F150LP UV emission (left panel) and the 12-17Ā m slice from MIRI/MRS Channel 3 (right panel). The UV emission (map and grey contours) clearly shows the six young, massive and compact āsuper star clustersā (SSC1, SSC2, SSC3, SSC4, SSC5, SSC6) revealed in this galaxy (Thuan etĀ al. 1997). SSCs1-6 clusters have ages and masses in the range Ā Myr and Ā Mā (Reines etĀ al. 2008; Adamo etĀ al. 2010; see also Tab.Ā 2), and luminosity of Ā Lā, all within āĀ Ā pc (Thuan etĀ al. 1997). One of them, SSC1, is also known for its enhanced obscuration (Hunt etĀ al. 2001). The HeĀ IIĀ 4686 high-ionization emission revealed with MUSE (white dotted contours; seeing ā) follows the UV emission and peaks around SSCs4,5. The HST H black dashed contours highlight the three additional actively young star-forming regions, S3-Pa, S7 and S8, discovered by Thompson etĀ al. (2009) by analyzing SBSĀ 0335-052Ā E Pa emission. The MIRI/MRS image (right panel) reveals the clear MIRI/MRS PSF pattern due to a very bright point source located at the position of SSC1 as well as extended emission covering the other stellar clusters, as we will discuss in Sec.Ā 3.
3 Data Analysis and Methods
By a first visual inspection of the StageĀ 3 products, we noticed a point-source like emission dominating the continuum and emission-line fluxes close to the position of SSC1 (see Fig.Ā 1 right panel). To spatially investigate the source of emission (Sec.Ā 4 and Sec.Ā 5), it was essential to determine the spatial extent of the underlying extended emission. This was achieved by accurately modeling and isolating the point source contribution, ending up with a point-source subtracted dataset and a point-source spectrum. The details of this subtraction procedure are described in App.Ā A. In Sec.Ā 3.1 and Sec.Ā 3.2, we describe the data analysis applied to the point source and extended emission spectra.
3.1 Emission-line fitting
In each spectrum we fitted several MIR emission lines, performing the line fitting locally around each line in a km/s window centered at the observed wavelength of the line of interest. In particular, we used a combination of a Gaussian profile and a one-degree polynomial to account for emission-line and surrounding continuum emission in each spaxel of the point-source subtracted band datacube slices, as well as in the 1-D spectrum of the bright point-source and the spectrum extracted at the ULX position (shown in Fig.Ā 1). The 1-D spectrum of the bright point source is obtained by the best-fit model of the PSF photometry analysis described in App.Ā A, while the ULX spectrum is extracted from a region with variable diameter depending on the channel, matching the FWHM of the PSF (i.e., Ā 0.4ā, 0.5ā, 0.6ā, 0.9ā (Law etĀ al. 2023). Table 1 compiles the list of emission lines analyzed, together with their rest-frame wavelengths and ionization potentials (I.P.).
Line | I.P. | |
(m) | (eV) | |
[FeĀ II] | 5.34 | 7.90 |
[ArĀ II] | 6.98 | 15.76 |
Pf | 7.46 | 13.6 |
[NeĀ VI] | 7.65 | 126.25 |
[ArĀ III] | 8.99 | 27.63 |
[SĀ IV] | 10.51 | 34.79 |
Hu | 12.37 | 13.6 |
[NeĀ II] | 12.81 | 21.56 |
[NeĀ V] | 14.32 | 97.12 |
[NeĀ III] | 15.56 | 40.96 |
[SĀ III] | 18.71 | 23.34 |
[NeĀ V] | 24.32 | 97.12 |
[OĀ IV] | 25.89 | 54.95 |
Fig.Ā 2, 3 and 4 show the point source spectrum (SSC1), the spectrum extracted from the spaxels covering the ULX position, and a spectrum from the north-west region covering the clusters SSCs3,4,5,6 and S7 and the peak of UV emission (see Fig.Ā 1), respectively, as well as the best-fit (in red) for the MIR emission lines. In the following, we consider line detections as those with fluxes with signal-to-noise (S/N) higher than 3 and upper limits for measurements with S/N2. In particular, we note that emission lines typically observed in AGN (e.g., Dasyra etĀ al. 2024), such as [NeĀ VI]Ā Ā 7.65 (tentatively fitted only in the bright point source spectrum with ), [MgĀ V]Ā Ā 5.61 (I.P.Ā Ā eV) and [MgĀ VII]Ā Ā 5.5 (I.P.Ā Ā eV) are not detected.
3.2 Dust attenuation
The emission-line fluxes were converted from the original MJy/sr into erg/s/cm2/arcsec2 and corrected for the total foreground Galactic reddening along the line of sight using the values determined for this object by Berg etĀ al. (2022) and the Gordon etĀ al. (2023) dust extinction relationship.
When exploring the spectrum of the point source SSC1, we noticed a wide absorption in the 7.5-12.5m range, corresponding to the silicate feature at 9.7Ā m (highlighted in grey in Fig.Ā 2, 3 and 4). This feature (revealed also in SBSĀ 0335-052Ā E Spitzer spectrum; Houck etĀ al. 2004) pinpoints the presence of dust within this region, as already reported by Hunt etĀ al. (2014). Thus, an intrinsic dust attenuation correction has to be applied to the line fluxes as well. Before using the continuum to determine the level of dust attenuation, there are few extra corrections we needed to implement to the datacubes. First, we limited the impact of fringing on the continuum spectra by applying an additional 1D fringing correction using the JWST pipeline tools. Also, the data exhibit discontinuities, or ājumpsā, in the flux normalization among the different bands, which we corrected by fixing the normalization to Band A in Channel 1. With the continuum corrected and optimized for the analysis, we used the Continuum and Feature Extraction software (CAFE, Marshall etĀ al. 2007; Diaz-Santos et al., in prep.) to model the continuum of both the point-source spectrum and the point-source subtracted (extended) continuum in the range Ā m from Channel 1, 2 and 3 band datacubes (see App.Ā A). CAFE uses a combination of blackbody emission at different temperatures together with a custom attenuation law (compatible with Gordon etĀ al. 2023) to describe the overall shape of the continuum at near-to-mid infrared wavelengths. For the case of SBS 0335-052Ā E only two dust components were considered: a cool component, whose temperature falls in the range 40-100 K, and a warm component, with temperature in the range 100-500 K. While both spectra, the point-like source and the extended emission, needed a cool component of 80 K, only the point-like source required an additional warm component with 150 K to model the continuum shape. A more detailed description of the dust attenuation correction will be presented in del Valle-Espinosa et al., in prep. As a result, the peak of dust attenuation corresponds to the point source SSC1 with , while for the extended region we considered an average value of . We used the derived values and the Gordon etĀ al. (2023) dust extinction relationship to correct for intrinsic dust attenuation.
Although the extinction curve at the MIR wavelengths is almost flat, lines located in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, 3 and 4) can suffer from large attenuation. In the case of SBSĀ 0335-052Ā E, the dust-corrected fluxes of almost every emission line are increased by a factor of () in the point-like source (extended) spectrum, except for the [ArĀ III]Ā and [SĀ IV]Ā lines ā which fall within the silicate feature ā whose correction factor is (). This also implies that only emission-line ratios including [ArĀ III]Ā and [SĀ IV]Ā lines can be affected by dust attenuation.
3.3 PSF and pixel size matching
In this paper, we want to diagnose the main ionization source by comparing emission-line ratios and models. Since MIRI/MRS channels have a wavelength-dependent PSF and a different pixel size, we needed to convolve and rescale our emission-line maps before calculating the ratios of lines in different regions of the MIR spectrum. We avoided taking into account the complexity of the PSF profile in each sub-channel and wavelength as we did for the PSF subtraction procedure since it would have added further artefacts and uncertainties. Instead, we considered the linear relation between the MIRI/MRS PSF FWHM and found by Law etĀ al. (2023; see their Eq.Ā 1). Then, we convolved the line maps with a 2-D Gaussian kernel set to a value equal to the difference in quadrature of the PSF FWHMs at the wavelengths taken into account, similarly to Jones etĀ al. (2024). Specifically, we created two sets of emission-line maps, convolving them either to the [NeĀ III]Ā Ā 15.56 or the [OĀ IV]Ā Ā 25.89 PSFs (Ā px, 4.5Ā px, ā, 0.9ā, Ā pc, respectively). Finally, we rescaled all the emission maps to ChannelĀ 3 pixel size (0.245ā), where the majority of the emission lines taken into account in this work are located.
3.4 Photoionization and shock models
In order to understand the main source of ionization in Sec.Ā 4 and Sec.Ā 5, we compared the observed emission-line ratios to different sets of photoionization and shock models from the literature. In particular, we considered:
ā¢
Simple stellar population (SSP) models from MartĆnez-Paredes etĀ al. (2023; MP23 hereafter)111Available in the Mexican Million Models Database: https://sites.google.com/site/mexicanmillionmodels/; http://3mdb.astro.unam.mx:3686/;
ā¢
SSP models with and without the ionizing output from X-ray binaries (XRBs) from Garofali etĀ al. (2024; G24 hereafter);
ā¢
Intermediate-massive black hole (IMBH) models from Richardson etĀ al. (2022; R22 hereafter);
ā¢
Shock models from Flury etĀ al. (2024; F24 hereafter) and the Mexican Million Models Shock Database222http://3mdb.astro.unam.mx:3686/, described in Alarie & Morisset (2019; AM19 hereafter).
For all these models, we considered only the two lowest metallicity grids available (i.e., 5% and 10% Zā models for G24, R22, F24, and 2% and 10% Zā for MP23), in accordance with SBSā0335-052Ā Eās metallicity range (Papaderos etĀ al. 2006; Nakajima etĀ al. 2024). Also, we took into account only bursts of star-formation with ages below 100Ā Myr (upper limit on age of SBS 0335-052Ā E) to match SBS 0335-052Ā E properties (Izotov etĀ al. 1997; Wofford etĀ al. 2021; Berg etĀ al. 2022; Mingozzi etĀ al. 2024, Martinez et al. in prep.). For all the grids, we consider the ionization parameter log() between -4 and -1.5 (apart from R22 where log() reaches -0.5). For MP23, R22 and shock models, we could also test the effect of different densities (Ā Ā cm-3 in MP23, R22; Ā Ā cm-3 in shock models), finding small differences that only slightly impact the results shown in Sec.Ā 4, as we discuss in Sec.Ā 5. Here below we give a summary of the characteristics of all these models.
MP23 SSP models combine CLOUDY (Ferland etĀ al. 2017) with the revised version of the Bruzual & Charlot (2003) stellar population synthesis models introduced in Plat etĀ al. (2019; C&B models), spanning across several parameters (see their TableĀ 4), including a Kroupa (2001) and top-heavy (x030) initial mass functions (IMFs) with a mass range up to 100 and 300 Mā. C&B models accurately cover the evolution of O and B stars, including the Wolf-Rayet (WR) phase, whose contributions last a few Myr. C&B models also include hot post-Asymptotic Giant Branch (pAGB) stars, also known as hot low-mass evolved stars (HOLMES; StasiÅska etĀ al. 2008), which emit a strong UV continuum ionizing and exciting the surrounding medium at ages between 70 and 500 Myr (peak around 100 Myr).
G24 modeled simple stellar populations (SSPs) and a population of ultraluminous X-ray sources (ULXs), combining the ionizing impact of a SSP with XRBs, creating āsimple X-ray populationsā (SXPs). As described in detail in G24, XRBs form in multiple generations following a burst of SF (Ā Myr), thus can produce ionizing photons on longer timescales than single massive stars (up to Ā Myr). Also, when present, they dominate the X-ray power output of star-forming galaxies and have a luminosity that scales with SFR. G24 used binary population synthesis models from Fragos etĀ al. (2013) and burst ages ranging from 1 to 20Ā Myr. For the SSPs, they used the stellar population synthesis code BPASS v2.2 (Eldridge etĀ al. 2017). Then, they created the SSP and SXP models with CLOUDY considering different values of gas-phase metallicities and ionization parameters, and a close geometry. Concerning the chemical composition of the CLOUDY models, G24 followed the prescriptions used in R22.
R22 made emission line predictions for Ā Mā IMBHs with CLOUDY, taking into account two extreme spectral energy distributions (SED) - a ādisk-plawā and a āqsosedā - for the incident radiation field (see R22 Sec.Ā 2.1) - and a starburst continuum including WR contribution, using the stellar population synthesis code BPASSĀ v2.0 (Stanway etĀ al. 2016) with a Kroupa IMF. To address different physical uncertainties, R22 models at different values of metallicities and ionization parameters include both coincident and non-coincident mixing of the IMBH and starlight excitation, and open and closed geometries. In particular, in this paper we show the 20Ā Myr instantaneous burst of star formation grids (in SSPs [OĀ III]Ā 5007/HĀ peaks at 20Ā Myr; Xiao etĀ al. 2018), with coincident mixing and closed geometry. The line ratios do not vary significantly in the non-coincident outputs, but can be slightly affected by the choice of geometry, as we comment in Sec.Ā 4.
AM19 shock models are made with Mappings-V (Sutherland & Dopita 2017), a plane-parallel geometry and the same prescriptions of the well-known Allen etĀ al. (2008) models, but spanning a broad range of gas-phase metallicities and densities. F24 models are made similarly with Mappings-V, but uniformly implemented empirical abundance patterns and dust depletion treatment as well as updated atomic data. Generally, two sets of models are taken into account: pure shocks and shocks+precursor. In this work, we consider only shocks+precursor models, since they also consider the gas entering the shock front. The grids we show span values of shock velocity (Ā km/s) and the magnetic field parameter (Ā Gauss), which are the main parameters regulating the shock ionization spectrum and the effective ionization parameter. We find that AM19 and F24 shock models behave similarly in the range of parameters considered, thus in Sec.Ā 4 we show only F24 grids.
4 Results
4.1 Emission-line maps
Fig.Ā 5 shows the emission line maps of the extended MIR emission lines obtained after the point source subtraction: PfĀ Ā 7.49m, HuĀ Ā 12.37m, [NeĀ II]Ā Ā 12.81m, [SĀ III]Ā Ā 18.71m, [ArĀ III]Ā Ā 8.99m, [SĀ IV]Ā Ā 10.51m, [NeĀ III]Ā Ā 15.56m, [OĀ IV]Ā Ā 25.89m, and [NeĀ V]Ā Ā 14.32m, in order of ionization potential (see Tab.Ā 1). The maps of all the lines but [SĀ III]Ā and [OĀ IV]Ā (Channel 4; ā, Ā pc) are convolved with the spatial resolution at the [NeĀ III]Ā wavelength (ā, Ā pc) and rescaled to Channel 3 pixel scale (0.245ā). The FOV varies from 3.2ā3.7ā, to 4.0ā4.8ā, 5.2ā6.2ā and 6.6ā7.7ā for Channel 1 (Pf), 2 ([ArĀ III], [SĀ IV]), 3 (Hu, [NeĀ II], [NeĀ V], [NeĀ III]) and 4 ([SĀ III], [OĀ IV]), respectively. In the [SĀ III]Ā map panel we indicate the positions of SBSĀ 0335-052Ā E several star-forming regions as we did in Fig.Ā 1.
PfĀ and HuĀ maps follow the HĀ distribution, which peaks in young star-forming regions (Ā Myr), such as the position of SSC1 (point source) and SSC2 as well as S7 and S3-Pa, slightly shifted with respect to the other SSCs and very bright in Pa (Thompson etĀ al. 2009). [NeĀ III]Ā and [SĀ IV]Ā have a very similar distribution with their emission extending across the MIRI/MRS FOV, due to their similar I.P. (see Tab.Ā 1). Higher ionization emission such as [NeĀ V]Ā and [OĀ IV]Ā mainly comes from the north-west portion of the galaxy, covering the older clusters SSCs3-4-5-6 (Ā Myr) and peaks around the S7 position. The UV emission (HST ACS 150LP; Fig.Ā 1) shows a different pattern with respect to MIR and optical hydrogen lines, tracing the six SSCs and peaking in the [NeĀ V]Ā and [OĀ IV]Ā high ionization region. We highlight that HeĀ IIĀ 4686 (similar I.P. to [OĀ IV]) emission studied by Kehrig etĀ al. (2018) with MUSE data closely follows the UV contours.
We notice that the peak of [NeĀ III]Ā and [SĀ IV]Ā emission is close to the ULX position revealed by Prestwich etĀ al. (2013; see also ) with Chandra (magenta circle). However, there is no clear correspondence between the ULX X-ray emission and high ionization lines, since the [NeĀ V], [OĀ IV]Ā and UV emission peaks are shifted towards S7. We highlight that the ULX extracted spectrum is also characterized by [FeĀ II]Ā 5.34m emission, that is not shown in Fig.Ā 5 since it does not show clear extended emission (see Fig.Ā 12).
The region where we subtracted the point-source emission, located at the position of SSC1, shows either a reduced or tentative emission (grey regions with ) in the line maps. This can be clearly seen in the HuĀ and [NeĀ V]Ā maps. This could be due to a real lack of extended emission around SSC1 or to uncertainties in the PSF subtraction procedure. As shown in Fig.Ā 2 the point source SSC1 shows almost all the emission lines reported in Fig.Ā 5, apart from [NeĀ V]Ā and [ArĀ II]. The point source spectrum also shows clear [FeĀ II]Ā emission (see also Fig.Ā 12).
4.2 Emission-line ratio maps
Fig.Ā 6 shows [NeĀ III]/[NeĀ II], [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ II]Ā line ratio extended emission maps with a color-coding chosen to enhance the peak regions to aid with the interpretation of the diagnostic diagrams. [NeĀ III]/[NeĀ II]Ā is known to be a good tracer of the ionization parameter log(), while [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā are found to be sensitive also to the hardness of the ionizing spectrum (Richardson etĀ al. 2022; Garofali etĀ al. 2024). Indeed, these line ratios do not show the same behavior. The peak of [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā maps are located around region S7, aligning with the peak in HĀ emission (see black contours). While [NeĀ III]/[NeĀ II]Ā is also enhanced in S7, it instead peaks at the position of S3-Pa, also bright in HĀ (see black contours). SSC3, SSC4, SSC5 and SSC6 are characterized by high ionization emission, but lower [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios than S7. As mentioned above, the extended region covering SSC1 (point source) and SSC2 does not have sufficient S/N in [NeĀ II], [NeĀ V]Ā and [OĀ IV]. Finally, S8 does not have either [NeĀ V]Ā or [OĀ IV]Ā emission but shows intermediate values of [NeĀ III]/[NeĀ II].
4.3 Ionized gas kinematics maps
Fig.Ā 7 shows the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā velocity and intrinsic velocity dispersion maps. The velocity maps are centred on the systemic velocity of this galaxy according to previous optical studies (Ā km/s; Moiseev etĀ al. 2010)333 We centred the [NeĀ III]Ā velocity map -24Ā km/s below the systemic velocity Ā km/s found by Moiseev etĀ al. (2010). In particular, we find this specific systematic shift in HuĀ and [NeĀ II]Ā (Ch.Ā 3A), and [NeĀ III]Ā (Ch.Ā 3C) velocity maps also obtained for the Stage 3 products, so it does not depend on the point-source subtraction procedure explained in Sec.Ā A. We suspect that this shift is not connected to the physical conditions of the gas, as it corresponds to Ā half the wavelength step in ChannelĀ 3. Also, Pf, [ArĀ III]Ā and [OĀ IV], located in different channels and characterized by different I.P., show consistent velocity maps to [SĀ IV]Ā and [NeĀ V]., while the intrinsic velocity dispersion maps are obtained by subtracting in quadrature the instrumental broadening dependent on wavelength from the observed velocity dispersion (see Eq.Ā 1, Jones etĀ al. 2023).
The MIRI/MRS FOV is limited to the central Ā kpc2 of SBSĀ 0335-052Ā E and does not cover the entire ionized gas distribution traced in HĀ by previous studies (e.g., Moiseev etĀ al. 2010; Herenz etĀ al. 2023), with signs of outflowing gas extending up to Ā kpc (Herenz etĀ al. 2023; in the north-west direction, along the galaxy minor axis). Thus, despite Moiseev etĀ al. (2010) showing that on larger scales there is a disk-like north-west to south-east gradient, we only see the inner regions known to be characterized by perturbed kinematics (Izotov etĀ al. 2006; Moiseev etĀ al. 2010; Herenz etĀ al. 2017; 2023). In particular, [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā maps trace progressively higher ionization gas (Tab.Ā 1) and show a consistent velocity pattern, with a blue-shifted emission around the position of S3-Pa and in the north-west portion of the MIRI/MRS FOV (on the right of S7 and at the top of S8), and red-shifted emission from north to south and in the east region.
Despite the kinematic complexity shown in the velocity maps, the emission lines are well-fitted with a single Gaussian, broadening in the north-west region of the MIRI/MRS FOV. Interestingly, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā intrinsic velocity dispersion maps seem to peak in slightly shifted positions. They all show a curious āVā shape (less visible in [NeĀ V]) with enhanced that peaks in SSCs 4-5 in the [NeĀ V]Ā map, Ā px north than that in [NeĀ III]Ā and further north in [SĀ IV]. We notice that the region with enhanced in the [NeĀ V]Ā map is located close to the peak of the [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ II]Ā line ratios (Fig.Ā 6) tracing the hardening of the radiation field.
The regions where [NeĀ III]Ā and [SĀ IV]Ā peaks probably have no [NeĀ V]Ā emission due to their lower ionization (they also partly show [NeĀ II]Ā emission; Fig.Ā 5) and are characterized by enhanced red-shifted velocities, possibly showing a flow of gas moving away from the [NeĀ V]Ā high ionization region. This picture is consistent with what found by Izotov etĀ al. (2006) with high-resolution () VLT/GIRAFFE ARGUS spectra, showing double-peaked HĀ profiles in the region north of SSCs 4-5 (see their Fig.Ā 7a and 8). In particular, they interpreted these features with the presence of an expanding shell of ionized gas at Ā km/s, while the HĀ broadening in SSCs 4-5 with higher dynamic activity than the surroundings.
From a visual comparison of the maps, the 15Ā kpc outflow revealed by Herenz etĀ al. (2023) in HĀ is not clearly connected to the central kpc kinematics and the location of the sources S7, SSC4-5. Interestingly, previous optical integral-field spectroscopy studies have also revealed an offset between the expanding superbubble shells and the star clusters driving the global outflow in galaxies similar to SBSĀ 0335-052Ā E (e.g., Martin etĀ al. 2024). Finally, we stress that no ionized outflow is detected in SBSĀ 0335-052Ā E HST COS UV absorption lines in correspondence of the sources S7, SSC4-5 (Xu etĀ al. 2022; Parker etĀ al. 2024).
4.4 Diagnostic diagrams
In this paper, we explore the following diagnostic diagrams widely used in the literature (see Hao etĀ al. 2009; Weaver etĀ al. 2010; Inami etĀ al. 2013, but also R22, MP23, G24):
ā¢
[NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II], color-coded as a function of [OĀ IV]/[NeĀ III]Ā (Fig.Ā 8);
- ā¢
- ā¢
We highlight that [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II]Ā represents the most reliable diagnostic diagram explored here, since these lines lie in Channel 3, where the attenuation law is flat, thus they are minimally impacted by PSF variations and dust correction. [SĀ IV]/[NeĀ II]Ā instead can be affected by dust reddening, while [OĀ IV]/[NeĀ III]Ā is the ratio most affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing, given that [NeĀ III]Ā is located in ChannelĀ 3c and [OĀ IV]Ā in ChannelĀ 4c. For these reasons, we mainly focus this section on [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II], moving the other two diagnostic diagrams in App.Ā B.
[…]
Fig.Ā 8 (same for Fig.Ā 13 and Fig.Ā 14, App.Ā B) show four panels where we overlay the diagnostic diagrams with a different set of low-metallicity models (see Sec.Ā 3.4):
ā¢
MP23 2.5Ā Myr SSPs, with top-heavy IMFs and Ā Mā (upper left panel);
ā¢
G24 SSPs and SXPs grids at 5, 10 and 20Ā Myr (upper right; the hexagon represents the highest line ratio each grid can reproduce);
ā¢
R22 IMBHs grids with Ā Mā, qso SED and 20 Myr SSP, with closed geometry and 4, 8, 16 and 100% AGN fraction, highlighting that beyond 16% the SED is completely dominated by the IMBH (lower left panel);
ā¢
F24 shock models with precursor (lower right panel).
It should be noted that for the MP23 SSP grids, the low metallicity 2.5 and 100Ā Myr SSPs show the highest line ratios for young and old stars, respectively. These two grids are the only ones (partially) able to cover the x- and y- ranges shown. However, we discarded the 100Ā Myr grid because old stars struggle to explain the observed number of ionizing photons observed in this system as we discuss in Sec.Ā 5.2.1. Also, we notice that the choice of the IMF does not affect the line ratios, but the choice of Ā Mā strongly enhances [NeĀ V]/[NeĀ II]. All the other MP23 grids and G24 SSPs and SXPs grids below 5Ā Myr fall outside the shown x- and y-ranges. R22 IMBHs grids with Ā Mā are also not shown because they cannot reproduce log([NeĀ III]/[NeĀ II]) higher than . The grids with Ā Mā instead can overlap to the observed line ratios at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II].
Fig.Ā 8 (same for 13 and 14) scatter plots show the extended emission measured with our data (down-pointing arrows for upper limits), as indicated in the legend. We stress again that the point source (star symbol) has only tentative [NeĀ V]Ā with . The cyan point indicates the [NeĀ V]Ā and [NeĀ II]Ā upper limits and the [NeĀ III]Ā and [OĀ IV]Ā detections from Spitzer flux measurements (Houck etĀ al. 2004; Hao etĀ al. 2009), which cover the entire region covered by MIRI/MRS. In particular, Houck etĀ al. (2004) reported that the Spitzer image had a prominent diffraction ring and was thus indistinguishable from the image of a point source (possibly SSC1 as we reveal in our data).
From a quick look at Fig.Ā 8 (same for Fig.Ā 13, App.Ā B), all SBS 0335-052Ā E MIR line ratios lie outside of the ācanonicalā SF locus, ending in the āAGNā designated region according to Inami etĀ al. (2013) and Richardson etĀ al. (2022) criteria (see also e.g., Goold etĀ al. 2024). This is already suggested by Spitzer upper limits (cyan point), but with MIRI/MRS data we can tell much more about the conditions of the ionized gas on a spatially resolved basis. Clearly, the [NeĀ III]/[NeĀ II]Ā and [NeĀ V]/[NeĀ II]Ā diagnostic diagram disentangles increasing ionization parameter and radiation hardness sequences, with both [NeĀ III]/[NeĀ II]Ā and [NeĀ V]/[NeĀ II]Ā progressively increasing with the log() of different models (i.e., following the log() arrow shown in the various panels). [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā (color-coding), being more sensitive to the radiation hardness, increase also along the perpendicular direction of the increasing log() arrow, at decreasing [NeĀ III]/[NeĀ II]Ā values. This effect is reproduced by harder ionization in the models, such as lower gas metallicity or older star ionization in SSPs models, and higher AGN fractions in IMBH models.
Another important takeaway from the top two panels of Fig.Ā 8 (same for Fig.Ā 13 and 14, App.Ā B) is that all the low-metallicity young SSPs grids struggle to reproduce the observed MIR emission line ratios of SBSĀ 0335-052Ā E. This suggests the need for a further ionization source. However, SXP models, including the contribution of X-ray binaries, show similar limitations. The only models that can reproduce SBSĀ 0335-052Ā E MIR [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios are models including the contribution of an IMBH (%), spanning a range of log(). However, we highlight that these IMBH models still struggle to reproduce the highest [OĀ IV]/[NeĀ III]Ā line ratios (Fig.Ā 14). Literature BCDs [SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā observed with Spitzer follow the Groves etĀ al. 2008 empirical relation shown in Fig.Ā 13 (magenta line). This indicates that low-metallicity SSP, SXP and shock grids generally struggle to reproduce BCDsā lower ionization MIR lines than [NeĀ V].
Given the critical densities of the MIR lines, another property that can affect the parameter space covered by these line ratios is the gas density. For instance, [Ne V] 14.3 is part of a doublet, showing a second transition at 24.3 m, with lower critical density ( cm-3 at 24.3, 14.3 m). This makes increasing [Ne V] 14.3/[Ne V] 24.3 ratios a good electron density () diagnostic in the range  cm-3 (e.g., FernÔndez-Ontiveros et al. 2016). We do not detect [Ne V] 24.3, which could be either due to the worsening of fringing in Channel 4 or to a high density of the [Ne V] emitting region. We favor the latter explanation given the fact that the [Ne V] emitting region shows high density according to previous UV and optical studies (i.e., ([Fe III]) ; (C III]) ; ([Ar IV]) ; (N IV]) ; Mingozzi et al. 2022). Interestingly, only models with  cm-3 seem able to reproduce a [Ne V] 14.3 stronger than [Ne V] 24.3. However, these high-density grids are slightly shifted towards lower [Ne III]/[Ne II] and [O IV]/[Ne III] line ratios (dotted grids in all diagrams), and therefore tend to make the models less consistent with the data.
Overall, all these considerations suggest that none of the current state-of-the-art models can perfectly match our observed MIR line ratios. In Sec.Ā 5, we discuss the feasibility of these results, also comparing them with previous spatially resolved studies performed on this galaxy.
5 Discussion
As shown in the previous sections (see also Fig.Ā 1, right panel), the MIRI/MRS data of SBS 0335-052Ā E are characterized by a bright point source emission, clearly showing the instrument PSF structure. As explained in Sec.Ā 3 and App.Ā A, we were able to model the point source contribution and accurately separate it from the underlying extended emission. Tab.Ā 2 shows a summary of the information that MIRI/MRS can give us on the several star-forming regions of SBSĀ 0335-052Ā E, as well as information gathered from the many previous works in the literature on this intriguing object to help with the interpretation of our data and the models presented in SectionĀ 3.4. In the following subsections, we discuss the origin of the point source and extended emission, their main ionization sources and we connect our results with the high- universe.
Region | Age [Myr] | Mass [] | 12+log(O/H) | Notable features | |
SSC1 | 3.0 | 4.7 | 7.31 | Point source in MIRI-MRS and NICMOS 1.6Ā m (T09) | |
SSC2 | 3.0 | 3.7 | SSC1-2 uniquely has dust and H2, and thermal radio emission (J09) | ||
SSC3(+S3-Pa) | 7.0 | 7.1 | 7.31 | WR Blue bump (P06; K18) | |
SSC4 | 11.0 | 1.1 | 7.27 | SSCs4-5 characterized by ([NeĀ V]) peak and tentative radio emission (J09) | |
SSC5 | 13.0 | 2.9 | Point source in NICMOS 1.6Ā m (T09) | ||
SSC6 | 11.0 | 2.6 | ā | Point source in NICMOS 1.6Ā m (T09); ([SĀ IV]) and ([NeĀ III]) peak | |
S3-Pa | (?) | ā | 7.38 | Bright H lines (T09) | |
S7 | (?) | ā | 7.10 | Bright H lines (T09); [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā peak | |
S8 | ā | ā | 7.21 | ||
ULX | ā | ā | ā | Revealed by P13 (see also K18) | |
Region | [NeĀ III]/[NeĀ II] | [NeĀ V]/[NeĀ II] | [OĀ IV]/[NeĀ III] | [SĀ IV]/[NeĀ II] | [NeĀ V]Ā 14.3m |
SSC1 | |||||
SSC2 | |||||
SSC3+S3-Pa | ā | ||||
SSC4-5 | ā | ||||
SSC6 | ā | ||||
S7 | ā | ||||
S8 | ā | ā | |||
ULX | ā |
5.1 The bright point source, SSC1
As introduced in Sec.Ā 3, the point source is the source SSC1 (Fig.Ā 1). SSC1 appears as a point source in the MIRI/MRS data possibly because of its dusty nature, which makes it particularly bright in the MIR. As described in Sec.Ā 3.2, when analyzing the MIR continuum, we find a very high attenuation for the point source spectrum (). Indeed, as shown in Fig.Ā 2, this spectrum is characterized by the silicate feature at 9.7Ā m, which is much fainter in the other regions of the galaxy. This is consistent with the fact that Hunt etĀ al. (2014) revealed dust continuum emission uniquely coming from the region around SSC1 using ALMA data (see their Fig.Ā 1). This region is also uniquely characterized by emission, tracing warm molecular gas, as first revealed by NICMOS data (Reines etĀ al. 2008; Thompson etĀ al. 2009). We confirm this with our data, observing several H2 rotational transitions around SSC1 (from S(1) to S(7)), which will be modeled and discussed in detail in a forthcoming paper (del Valle-Espinosa et al., in prep.). However, the emission may come from a more extended region, given that the H2 spatial distributions do not show the PSF pattern (see also Thompson etĀ al. 2009; their Fig.Ā 3). The point source region and close surroundings also show high HĀ and Pa equivalent widths (Reines etĀ al. 2008) and purely thermal radio free-free emission, which has been interpreted as due to active star-formation in a dense environment (i.e., Ā O7.5Ā V stars, Ā cm-3, 1.3Ā Mā/yr-1; Johnson etĀ al. 2009).
It should be noted that SSC1 also displayed typical PSF Airy ring in HST Near Infrared Camera and Multi-Object Spectrometer (NICMOS) NIR data, tracing Pa and 1.6Ā m continuum emission (0.2ā/px, comparable to MIRI/MRS) in Reines etĀ al. (2008); Thompson etĀ al. (2009). In particular, Thompson etĀ al. (2009) showed that high-resolution (0.075Ā ā/px) NICMOS camera 2 data could resolve the SSC1 emission and estimated the point source emitting region to be Ā pc. Thompson etĀ al. (2009) showed that also SSC5 and SSC6 show a point source structure in the NIR NICMOS 1.6Ā m continuum, with SSC5 seeming so compact (spatial extension of Ā pc) that could not be resolved even in their high-resolution NICMOS dataset. In our MIRI/MRS data, we do not reveal any point source emission corresponding to SSC5 and SSC6, implying that those regions of the galaxy are instead dominated by extended emission in the MIR.
All the elements listed above seem consistent with SSC1 being a very young embedded star-forming region (Hunt etĀ al. 2001; Reines etĀ al. 2008; Adamo etĀ al. 2010). Indeed, as shown in Fig.Ā 8 and Ā 14, a subset of 2.5Ā Myr SSP models from MP23 could tentatively explain the point source [NeĀ III]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios - despite the fact that they fall in the āAGNā regions of the diagnostic diagrams - and possibly the [NeĀ V]/[NeĀ II]Ā upper limit. However, SSP models, are not capable of reproducing the corresponding lower ionization [SĀ IV]/[NeĀ II]Ā line ratio (see Fig.Ā 13), which seems more consistent with R22 Ā Mā grids with an AGN fraction below 8%.
Interestingly, the point source spectrum is also characterized by [FeĀ II]Ā emission, mainly located around SSCs1-2. [FeĀ II]Ā is considered an important shock tracer, because of (i) its many levels with low excitation energies easily excited in shocked gas, (ii) its low ionization potential (Tab.Ā 1) that can make it ionized by FUV radiation from the shock front in the neutral H gas and (iii) is enhancement could be due to grain destruction by shocks (e.g., Koo etĀ al. 2016 and references therein). However, 5-10%Ā Zā shock models cannot reproduce the observed MIR line ratios.
Finally, as we will discuss in detail in Sec.Ā 5.2, Kehrig etĀ al. (2018) derived SBSĀ 0335-052Ā E HeĀ IIĀ 4686-ionizing budget (including at the position SSCs1-2; knotĀ B in their nomenclature), finding that it can only be produced by either single, rotating metal-free stars or a 0.05%Ā Zā binary population and a top-heavy IMF. Since HeĀ IIĀ and [OĀ IV]Ā have similar I.P., they are likely powered by the same ionizing source, which could also cause the lower ionization MIR emission. However, a 0.05%Ā Zā (or lower) metallicity is extremely low and does not match with the gas-metallicity measured in this object (i.e., %Ā Zā; Thuan etĀ al. 2004; Papaderos etĀ al. 2006; Izotov etĀ al. 2006), making this scenario unlikely.
Overall, there is not a straightforward explanation to justify all the characteristics of the point source SSC1 emission. Having said that, the models that can simultaneously explain the MIRI/MRS [OĀ IV]/[NeĀ III], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā are R22 IMBH models (see Fig.13Ā andĀ 14). While these grids are unable to predict the low [NeĀ V]/[NeĀ II]Ā upper limit, we cannot completely exclude that [NeĀ V]Ā upper limit is underestimated since the point source spectrum is affected by residual fringing in ChannelĀ 3.
5.2 Origin of the highly ionized extended emission
As shown in Fig.Ā 5, SBSĀ 0335-052Ā E is characterized by extended high ionization emission lines such as [OĀ IV]Ā 25.89 and [NeĀ V]Ā 14.3, localized mainly in the north-west region of the galaxy, covering SSCs 3-4-5-6 and S7 (all older than the point source SSC1; see Tab.Ā 2), with a peak of high ionization emission located in the proximity of S7. We highlight that [NeĀ V]Ā 14.3 has never been detected in BCDs before and SBSĀ 0335-052Ā Eās log([NeĀ V]/[NeĀ II])Ā values are higher than any upper limits proposed for BCDs (e.g., Hao etĀ al. 2009), probably because of the insufficient sensitivity of ISO and Spitzer. However, its optical counterpart (i.e., [NeĀ V]Ā 3426) has been previously detected in a subgroup of BCDs, including SBSĀ 0335-052Ā E (see Izotov etĀ al. 2004; Thuan & Izotov 2005; Izotov etĀ al. 2012; 2021; Berg etĀ al. 2021), which is the lowest metallicity object in which this line has been detected so far. [NeĀ V]Ā 3426 (as well as [NeĀ V]Ā 14.3,24.3) is usually considered a convincing proof of AGN activity (e.g., Abel & Satyapal 2008; Mignoli etĀ al. 2013), powered by the hard non-thermal AGN radiation, but has also been generally justified with radiative shock models in BCDs, using Allen etĀ al. (2008) models at solar and SMC (i.e., 20%Ā Zā) metallicities (e.g., Izotov etĀ al. 2012; 2021). However, as shown in Fig.Ā 8, 13 and 14, log([NeĀ V]/[NeĀ II])Ā and enhanced log([OĀ IV]/[NeĀ III]) cannot be explain by current state-of-the-art shock models at SBSĀ 0335-052Ā Eās metallicity (i.e., %Ā Zā).
The difficulty in accounting for SBSĀ 0335-052Ā Eās extended high ionization emission was first described by Herenz etĀ al. (2017), who showed a nebular HeĀ IIĀ 4686 map from MUSE that brightened towards the north-west at the rim of a starburst driven super-shell (see also Kehrig etĀ al. 2018), the HeĀ IIĀ shell that is out of our MIRI/MRS FOV. In the following sections, we discuss in detail the different scenarios which can explain the sources of the extended MIR emission, drawing insights from previous studies. In particular, it is important to consider that [OĀ IV]Ā has a similar I.P. to HeĀ IIĀ (Ā ryd), while [NeĀ V]Ā requires the presence of even harder radiation (Ā ryd), corresponding to the extreme ultraviolet and soft X-ray range.
5.2.1 Pure Bursts of Star Formation
MP23 and G24 SSP models considered in this work, fail to reproduce SBSĀ 0335-052Ā E [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II], [SĀ IV]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios, despite using different stellar libraries and IMFs (see Sec.Ā 3.4). This is not unexpected considering previous studies of this system.
As mentioned previously in Sec.Ā 5.1, Kehrig etĀ al. (2018) was only able to derive SBSĀ 0335-052Ā Eās HeĀ II-ionizing budget via either single, rotating metal-free massive stars (Ā MĀ Ā Mā stars, see their Sec.Ā 5.3.1) or a binary population with 0.05%Ā Zā metallicity and a top-heavy IMF (modeled with BPASS; see their Sec.Ā 5.3.2). However, a 0-0.05%Ā Zā metallicity is unphysically low for SBSĀ 0335-052Ā E (see also Wofford etĀ al. 2021), making this scenario unlikely. Kehrig etĀ al. (2018) discarded also Wolf-Rayet (WR) stars - massive stars (Ā Mā) with strong winds (e.g., Schaerer & Vacca 1998) - as the main source of ionization. Despite WR features being observed in this galaxy (in S3 and in the south-west region of the galaxy, below S8; Papaderos etĀ al. 2006; Izotov etĀ al. 2006; Kehrig etĀ al. 2018), the large number needed to justify the HeĀ II-ionization budget () would overcome by a factor of the SSCs total mass, depending on the IMF choice (Kehrig etĀ al. 2018). The same argument allows us to discard WRs as the main source of [OĀ IV]Ā and thus [NeĀ V]Ā emission.
Wofford etĀ al. (2021; see their SecĀ 3.1) proposed the presence of very massive stars (VMSs) - stars with masses above 100Ā Mā (Vink 2012) - from the blue-shifted [OĀ V]Ā 1371 detection in HST COS data, covering a 2.5ā circular region centered on the peak of the UV emission (see Fig.Ā 1). Thus, one might suspect that VMSs might represent the definite answer, hardening the radiation field and enabling highly ionized extended emission. However, VMSs, while hot and luminous, are known to have dense winds which become optically thick at the short wavelengths needed to power HeĀ IIĀ and [NeĀ V]Ā ionizing photons, as demonstrated by Sander (2022; see also Hawcroft et al. 2025, submitted). Although, a mechanism for the enhanced escape of the hard ionizing radiation from these winds has recently been proposed by Roy etĀ al. (2025), whereby the winds of massive stars are inhomogeneous due to intrinsic instabilities of the outflow. Roy etĀ al. (2025) propose that the higher the degree of inhomogeneity in the outflow, the higher the emergence of extreme UV fluxes. While this possibility has not yet been comprehensively evaluated, it may motivate revisions of stellar population models in the future.
Finally, a completely different scenario to explain the high ionization extended emission could be older ionizing stars, as suggested by previous works (e.g., Izotov etĀ al. 1997; 2006). MP23 low-metallicity SSPs models with ages of Ā Myr (upper limit on age of SBSĀ 0335-052Ā E; Izotov etĀ al. 1997; Papaderos etĀ al. 1998), would be dominated by post-AGB stars (specifically HOLMES) that could be capable of ionizing the gas with their strong UV continuum (StasiÅska etĀ al. 2008). Indeed, MP23 (see their Fig.Ā 2) showed that HOLMES have the peak production rate of H-ionizing photons exactly at 100Ā Myr. Interestingly, MP23 low-metallicity 100Ā Myr SSP models would perfectly lie on the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II], [SĀ IV]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratios shown in Fig.Ā 8,Ā 13,Ā 14. However, these models were disregarded because the maximum production of ionizing photons is Ā photonsĀ s-1 for a 1Ā Mā stellar population at SBSĀ 0335-052Ā E metallicity - implying that too large a population of HOLMES would be needed to match the ionization budget estimated by Kehrig etĀ al. (2018; i.e., Ā photonsĀ s-1). While these stars cannot be easily observed in a galaxy as distant as SBSĀ 0335-052Ā E, they are not revealed in large numbers even in nearby elliptical galaxies with a weak UV excess (e.g., Brown etĀ al. 2008), suggesting that they do not represent a significant source of UV emission as proposed.
Overall, we have discussed how state-of-the-art SSPs are not capable of reproducing high ionization extended emission observed in SBSĀ 0335-052Ā E. We also highlight that SSPs struggle to explain lower ionization MIR line ratios, such as [SĀ IV]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II]Ā (see Fig.Ā 13), even for values typically observed in BCDs with Spitzer (i.e., Groves etĀ al. 2008 relation in magenta). This implies either a general hindering problem in low metallicity massive stars in current stellar libraries or the real need for a further source of ionization in galaxies like SBSĀ 0335-052Ā E.
5.2.2 Ultra Luminous X-ray sources
A ULX is observed in SBS 0335-052Ā E (Prestwich etĀ al. 2013), as shown in Fig.Ā 1. In particular, ULX sources are X-ray binaries with Ā erg/s and the ULX fraction seems to anti-correlate with gas-phase metallicity (e.g., Mapelli etĀ al. 2010; 2011; Prestwich etĀ al. 2013). This could suggest that, given the characteristic low-metallicity gas of BCDs, ULXs could indeed play a role in their high-ionization emission.
The contribution of these X-ray sources were accurately included for the first time in the SXP grids of G24, which overlap with the range of [NeĀ V]/[NeĀ II]Ā ratios (including the values of the spectrum extracted at the ULX position, magenta diamond) observed in SBS 0335-052Ā E at the highest ionization parameter values (i.e., log()Ā ) and for stellar population ages between 5Ā Myr and 20Ā Myr (see Fig.Ā 8). However, the same grids over-predict the corresponding [NeĀ III]/[NeĀ II]Ā line ratios and, moreover, are not capable of reproducing log([OĀ IV]/[NeĀ III])Ā that we observe (Fig.Ā 14). We stress that G24 state-of-the-art model prescriptions could suffer from some inherent assumptions, such as the assumed theoretical scaling relationships linking X-ray luminosity with age and metallicity, while observationally there is significant scatter (see e.g., G24 Fig.Ā 1).
Another factor to consider is that the ULX location determined by Prestwich etĀ al. (2013) does not correspond to either the peak of the high ionization emission that we observe in the MIR or to the peak of the HeĀ IIĀ and UV emission. Indeed, Kehrig etĀ al. (2018) (see also Wofford etĀ al. (2021)) discarded the ULX as a possible source of HeĀ IIĀ ionization. In particular, analyzing Chandra data (see also Thuan etĀ al. 2004) they estimated the effective ionizing power at the peak of the X-ray emission and in other two regions of the galaxy, concluding that it is not enough to explain the observed HeĀ IIĀ emission. Accordingly, this implies that the ULX cannot explain [OĀ IV]Ā as well as [NeĀ V]Ā emission.
5.2.3 Shocks
Shocks models are generally considered an important source of ionization in MIR diagnostic diagrams and can be due to several phenomena, including cloud-cloud collisions, the expansion of HĀ IIĀ regions, outflows and/or supernovae (e.g., Groves etĀ al. 2004). Given the disturbed kinematic of this galaxy (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023), the presence of clear shells in the HĀ and HeĀ IIĀ morphology (Herenz etĀ al. 2017; Kehrig etĀ al. 2018) as well as diffuse non-thermal radio emission interpreted as an ensemble of compact supernova remnants expanding in a dense ISM (Hunt etĀ al. 2004), it is plausible that shocks are present and can play an important role. Unfortunately, the shell structure clearly visible in the HeĀ IIĀ morphology is not covered in our MIRI/MRS observations. This region is slightly north-west to the area where we see Ā km/s red-shifted velocities and intrinsic velocity dispersion enhancement in [SĀ IV]Ā and [NeĀ III]Ā (Fig.Ā 7).
Having said that, as shown in Figs.Ā 8, 13 and 14, low-metallicity F24 (and also AM19, not shown) shock models do not overlap with our data. The grids could match the range of observed [OĀ IV]/[NeĀ III], but cannot explain the corresponding [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios. Another hint towards shocks not being the primary source of ionization of the extended [OĀ IV]Ā and [NeĀ V]Ā emission is the complete lack of co-spatial [FeĀ II]Ā emission (see Sec.Ā 5.1). Finally, Kehrig etĀ al. (2018) also excluded shocks as the cause of the HeĀ IIĀ emission, given the lack of high ([OĀ I]Ā 6300) and lack of low [SĀ II]/HĀ line ratios, generally considered to be good shock indicators (e.g., Kewley etĀ al. 2019; Mingozzi etĀ al. 2024).
There might be other factors to consider when explaining the misalignment of shock models and extended high ionization emission. First, shock fronts have spatial scales much below our spatial resolution (down to Ā pc; Calzetti etĀ al. 2004), which could dilute shock-driven line ratios with other ionizing sources when integrating over spatial scales larger than a shock front. Second, it is possible that new shock models with an updated treatment of the radiative transfer with respect to MappingsĀ V could be in better agreement with our data (e.g., Godard etĀ al. 2024). However, these new models have not been tested yet at low-metallicity and currently take into account only velocities below 500Ā km/s. Overall, we cannot completely exclude the presence of shocks, but according to the available model predictions, they do not represent the main source of ionization of MIR lines in SBSĀ 0335-052Ā E.
5.2.4 The possibility of an IMBH
Recent works have suggested the presence of a Ā Mā IMBH in SBSĀ 0335-052Ā E, due to the detection of [NeĀ V]Ā 3426 (Hatano etĀ al. 2024), as well as possible variability in NEOWISE data (Ā m) over 12 years (Hatano etĀ al. 2023). Interestingly, the [NeĀ V]Ā 3426 was first observed by Thuan & Izotov 2005 and does not show any sign of variability (Hatano etĀ al. 2024). Hatano etĀ al. (2023) proposed also the presence of broad HĀ emission possibly due to the BH broad line region, from which they estimate a Ā Mā upper limit. However, this value is much larger than the mass of the different components of this galaxy (see Tab.Ā 2), raising doubts that this emission is tracing the BH broad line region (see also Sec.Ā 5.3), unless the system is not virialized (e.g., Bertemes etĀ al. 2025). As such, the existence of [NeĀ V]Ā 3426 and the observed NIR variability seem the two most feasible clues from the literature to suggest the presence of an IMBH in SBSĀ 0335-052Ā E.
In consideration of these suggestions of an IMBH, looking at Fig.Ā 8 and Fig.Ā 13, the models that uniquely cover the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios that we measure in SBSĀ 0335-052Ā E are the Ā Mā IMBH models from R22 with an AGN fraction 4-8-16% and log() between and (the closed geometry model can better reach the highest [NeĀ III]/[NeĀ II]). According to the R22 models, we can exclude the possibility of a BH with Ā Mā because this BH mass would imply too low log([NeĀ III]/[NeĀ II]) () and log([NeĀ V]/[NeĀ II]) () with respect to the observed data, regardless of the other parameter selection (i.e., AGN fraction, geometry, gas density, stellar age). We can also exclude R22 Ā Mā grids, which still cannot reach the log([NeĀ V]/[NeĀ II]) values that we observe (). R22 Ā Mā can overlap with the observed [NeĀ V]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II]Ā than what we observe.
A great advantage of the MIRI/MRS data is that we can spatially resolve line ratios and trace the ionization hardness (e.g., [NeĀ V]/[NeĀ II], [OĀ IV]/[NeĀ III]), finding their peak around/north SSCs 4,5 and S7 and with decreasing values towards south-east (see Fig.Ā 6). As already discussed in Sec.Ā 4.3, this peak region is characterized by the most disturbed kinematics with the presence of gas flows (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023). In particular, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā intrinsic velocity dispersion maps revealed an interesting āV shapeā in the north-west region, starting at the position of SSCs 4-5, high-resolution optical IFU data revealed doubled-peak HĀ and HeĀ IIĀ profiles, suggesting the presence of an expanding shell (Izotov etĀ al. 2006). In this scenario, SSC5 may correspond to the location of the IMBH. Another potentially relevant element is that SSC5 is also found to have a very compact (Ā pc) NIR emitting region not resolved in HST NICMOS (high resolution, 0.075Ā ā/px) data (Thompson etĀ al. 2009).
However, we do stress that even the R22 models do not perfectly overlap with our data, since they are not able to reproduce the combination of highest [OĀ IV]/[NeĀ III]Ā and [NeĀ III]/[NeĀ II]Ā values (see Fig.Ā 14). Many factors could be at play here with regards to the assumptions and parameters held within the models - each of which are difficult to constrain with observations alone due to the degeneracies between them. For instance, R22 predictions can vary as a function of the considered SEDs (R22 considered two extreme values, qso or disk-plaw) and assumptions on accretion rate and spin (incorporated as fixed parameters in the qso SED), as well as the choice of geometry (open vs closed). Also, we stress that [OĀ IV]/[NeĀ III]Ā can be affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing (see Sec.Ā 4.4).
5.3 Implications for high- galaxies and UV-optical diagnostics
An unexpectedly large population of objects with possible accreting black holes has started to be revealed at high- () with masses up to Ā Mā (e.g., Furtak etĀ al. 2023; Harikane etĀ al. 2023; LabbĆ© etĀ al. 2023; Scholtz etĀ al. 2023; Larson etĀ al. 2023; Greene etĀ al. 2024; Greene & Ho 2004; Maiolino etĀ al. 2023a; Kokorev etĀ al. 2024; Chisholm etĀ al. 2024; Hayes etĀ al. 2024; Cammelli etĀ al. 2025; as introduced in Sec.Ā 1). Not only has this led to questions concerning their rapid formation (in just 300Ā Myr) and their subsequent impact on galaxy evolution, but also opened up discussions on which methods can unambiguously identify them.
As previously mentioned, one of the most popular criteria to identify AGN activity so far has been via broad permitted emission lines (e.g., H). From a visual inspection of VLT/MUSE observations of SBS 0335-052 E MUSE, it appears that a broad H profile exists at both the position of cluster SSC1 (i.e., the MIR point source with no clear [Ne V] detection) and at the peak of He II, [Ne V], and UV emission, with FWHM up to  km/s (see Fig. 15 and App. C). In particular, both the H and [O III] 5007 emission lines require up to 3-4 Gaussian components to be well fitted, with also [O III] 5007 showing a fainter broad (up to FWHM  km/s) component. Interestingly, broad H and the need of components to reproduce the kinematics of the ionized gas has also been noticed in other nearby BCDs (e.g., James et al. 2009; del Valle-Espinosa et al. 2023; Komarova et al. 2021) and other high- analogs (e.g., Green Peas; Amor“in et al. 2024), with interpretations ranging from turbulent mixing layers, high-density gas, and radiatively driven winds. Despite the origin of this broadening still not being completely understood in these systems, there could be a connection with (some of) the broad-line galaxies revealed at high-.
Exploiting the use of high- analogs, Mingozzi etĀ al. (2024) used optical and UV spectra from the Cosmic Origins Spectrograph (COS) Legacy Archive Spectroscopic SurveY (CLASSY; Berg etĀ al. 2022; James etĀ al. 2022) to provide a toolkit to distinguish star formation and AGN activity at high-. In particular, CLASSY is a treasury of 45 nearby () star-forming galaxies, including SBSĀ 0335-052Ā E, generally characterized by lower stellar masses, higher SFRs and more extreme ionization fields than objects, as typically observed in the EoR. Specifically, Mingozzi etĀ al. (2024) proposed the CĀ III]Ā 1907,9/HeĀ IIĀ 1640 versus OĀ III]Ā 1666/HeĀ IIĀ 1640 as the most reliable UV diagnostic diagram (their Fig.Ā 6), being capable of separating the SF, AGN and shocks at low-metallicity. We re-propose this diagnostic diagram in Fig.Ā 9, showing the CLASSY sample compared with R22 models with the same parameters presented in FigureĀ 8 and discussed in SectionĀ 5.2.4. The fact that SBSĀ 0335-052Ā E (light-pink square) lies very close to the line separators was interpreted by Mingozzi etĀ al. (2024) as being due a combination of its low-metallicity environment and hard stellar radiation field. However, the R22 Ā Mā IMBH model with 8%Ā AGN fraction at the lowest metallicity and highest log() would also explain the UV line ratios observed in SBSĀ 0335-052Ā E. Since the UV emission line fluxes are from spectra integrated over the COS aperture (2.5ā; see Fig.Ā 1) covering the [NeĀ V]Ā 14.3 emitting region, the fact that the UV line ratios can be reproduced by the same set of models that explain part of the spatially resolved MIR line ratios indeed strengthens the validity of the R22 models in describing this object. This finding also highlights a previously unknown degeneracy of this diagnostic diagram in distinguishing between pure star-forming objects and objects with an AGN fraction Ā % (when considering R22 models).
The MIR wavelength range is not observable in high- targets, but JWST does provide coverage of [NeĀ V]Ā in the optical wavelength range (e.g., Cleri etĀ al. 2023; Scholtz etĀ al. 2023; Chisholm etĀ al. 2024). As mentioned at the beginning of Sec.Ā 5, optical [NeĀ V]Ā has been revealed in a few nearby BCDs, including SBSĀ 0335-052Ā E. Thus, it is interesting to see how SBSĀ 0335-052Ā E compares to other BCDs and high- targets within which this emission has been observed. In Fig.Ā 10 we show the diagnostic diagram HeĀ IIĀ 4686/HĀ vs [NeĀ V]Ā /[NeĀ III]Ā 3868 proposed by Chisholm etĀ al. (2024) to identify the presence of accreting IMBHs. Indeed, similarly to the MIR line ratios explored in Fig.Ā 8, both [NeĀ V]/[NeĀ III]Ā and HeĀ II/HĀ are relatively insensitive to metallicity and instead trace the ionization parameter and the radiation hardness (Shirazi & Brinchmann 2012; Cleri etĀ al. 2023; Chisholm etĀ al. 2024). Consistently with the MIR (Fig.Ā 8) and UV line ratios (Fig.Ā 9), SBSĀ 0335-052Ā Eās optical line ratios are also consistent with the R22 Ā Mā IMBH models with 8% AGN fraction. Interestingly, SBSĀ 0335-052Ā Eās [NeĀ V]/[NeĀ III]Ā flux ratio (pink square) is below the median values found for the 13 currently known nearby BCD [NeĀ V]Ā emitters (Izotov etĀ al. 2004; 2012; 2021; Thuan & Izotov 2005) and is of the value found for the galaxy studied by Chisholm etĀ al. (2024).
Upon reflection of our comparison with UV and optical emission line diagnostics, our analysis suggests that the MIR high-ionization emission that we observe in SBSĀ 0335-052Ā E could indeed be powered by an accreting IMBH. According to the best-fitting R22 models, the mass of the IMBH would be Ā Mā, placing it slightly above the scaling relation between and the host galaxy stellar mass, as seen for many objects at high- (e.g., Maiolino etĀ al. 2023a). However, given the fact that there are several open questions regarding the predicted high-ionization emission of other models, alternative sources of ionization cannot be ruled out. In particular, it is not clear if we have completely understood the atmosphere of low metallicity massive stars (see Sec.Ā 5.2.1), or how to constrain the different parameters within X-ray binaries, shocks and IMBHs models (see Sec.Ā 5.2.2, Sec.Ā 5.2.3, Sec.Ā 5.2.4). As such, our model-based conclusions can only be as firm as the models are realistic.
6 Conclusions
In this paper, we reported the first JWST MIRI/MRS data of a BCD, focusing on the galaxy SBSĀ 0335-052Ā E. This object is well-known for its extremely metal-poor environment, complex kinematics, high-ionization optical and UV emission lines, and the possible presence of an IMBH, which makes it the perfect analog of high- systems in the EoR. In particular, we used MIRI/MRS data to spatially resolve the MIR emission in this BCD (down to ā, Ā pc) through emission lines tracing different levels of ionization (e.g., [NeĀ II], [SĀ IV], [NeĀ III], [OĀ IV], [NeĀ V]) of the ionized gas. In this way, we can reveal the ionization structure of this galaxy and investigate the possible accreting IMBH or other non-stellar ionizing sources, building on the many previous works carried out since SBSĀ 0335-052Ā E discovery (Izotov etĀ al. 1990). In the following, we summarize our main findings:
ā¢
We revealed the presence of a point-source-like emission dominating the continuum and emission-line fluxes close to the position of one of the youngest and most embedded stellar clusters, SSC1 (Ā Myr, ; see Fig.Ā 1). We accurately modeled and isolated its emission (see Fig.Ā 2; Sec.Ā A), to study in detail its properties as well as the underlying extended emission. The point-source spectrum does not show very high-ionization emission (i.e., [NeĀ V]Ā 14.32), and is characterized by the strongest silicate feature at 9.7Ā m, H2 molecular gas emission and low ionization lines (i.e., [FeĀ II]Ā 5.34). As we discuss in Sec.Ā 5.1, there is not a straightforward explanation to explain all its characteristics. R22 Ā Mā IMBH models can well reproduce the detected MIR lines, but cannot explain the low [NeĀ V]/[NeĀ II]Ā upper limit.
ā¢
We revealed extended high ionization emission lines such as [OĀ IV]Ā 25.89 and [NeĀ V]Ā 14.32, localized mainly in the north-west region of the galaxy, covering SSCs 3-4-5-6 and S7 (all older than the point source SSC1). In particular, this is the first [NeĀ V]Ā 14.32 detection in a BCD (Sec.Ā 4.1, Fig.Ā 5).
ā¢
We spatially resolved line ratios tracing the ionization hardness (e.g., [NeĀ V]/[NeĀ II], [OĀ IV]/[NeĀ III]), finding their peak around/north SSCs 4,5 and S7 and with decreasing values towards south-east (Sec.Ā 4.2, Fig.Ā 6). Interestingly, we found log([NeĀ V]/[NeĀ II])Ā , higher than any upper limit proposed for BCDs with previous IR telescopes (Sec.Ā 4.2, Fig.Ā 6).
- ā¢
ā¢
We explored MIR diagnostic diagrams, comparing MIR line ratios with a set of state-of-the-art photoionization and shock models, finding that SBSĀ 0335-052Ā E lies outside of the ācanonicalā SF locus and can be best reproduced by R22 Ā cm-3 Ā Mā IMBH models with AGN fraction % (Sec.Ā 4.4; Fig.Ā 8, Fig.Ā 13). However, these models struggle to reproduce the highest [OĀ IV]/[NeĀ III]Ā MIR line ratios (Fig.Ā 14).
ā¢
The R22 models are consistent also with UV and optical emission-line diagnostics (Sec.Ā 5.3; Fig.Ā 9, Fig.Ā 10), strengthening the possibility of the presence of an accreting IMBH. On the other hand, star-forming models (regardless of including X-ray binaries) and shocks struggle to reproduce even MIR line ratios tracing lower ionization than [OĀ IV]Ā and [NeĀ V]Ā (Fig.Ā 13), typically observed in BCDs, suggesting possible model limitations. In Sec.Ā Sec.Ā 5.2 (see also 5.3) we also highlight how there are several open questions regarding the predicted high-ionization emission of other models, implying that alternative sources of ionization (e.g., very massive stars, Sec.Ā 5.2.1; Ultra Luminous X-ray sources, Sec.Ā 5.2.2; shocks, Sec.Ā 5.2.3) cannot be completely ruled out.
Overall, until models are adapted to fully replicate the ionizing sources, one way to step-forward in our understanding of the role that objects such as SBSĀ 0335-052Ā E have in interpreting the high- universe would be to perform detailed multi-wavelength modelling, i.e. utilizing the full X-rayāsub-mm information as constraints. Finding more near and far [NeĀ V]Ā emitters, as well as investigating more dwarf AGN candidates (e.g., Polimera etĀ al. 2022; Mezcua & DomĆnguez SĆ”nchez 2024; Wasleske & Baldassare 2024), covering a diverse range of properties, will also help to better understand the conditions required to produce this high-ionization emission.
Appendix A Point source subtraction procedure
We modelled the PSF for each MIRI/MRS band using vĀ 1.4.0 WebbPSF (Perrin etĀ al. 2014) and the function calc_datacube444https://webbpsf.readthedocs.io/en/latest/jwst_ifu_datacubes.html, taking into account the same native pixel scale and spectral binning of the observed data. We created the pipeline Stage 3 products by band, selecting the option ifualign, to have them aligned with the WebbPSF models. Then, we extracted m (m for Channel 4) slices around the emission lines taken into account in this paper (see Sec.Ā 3.1) from the Stage 3 band datacubes. We localized the point source position in each slice, modeling a 2D Gaussian distribution in a region of the spectrum with solely continuum emission, where the point-source structure is more enhanced (emission lines have also extended emission). Then, using photutils PSFphotometry we found the best-fit model of the point-source emission in each slice, wavelength-by-wavelength, putting very tight constraints on its position and allowing the PSF model normalization to vary to match the flux in a region times the PSF FWHM. According to WebbPSF and JWST documentation, the PSF FWHM in Channel 1, 2, 3 and 4 is (i.e., 2, 2.5, 3, 4.5Ā px), respectively.
FigĀ 11 displays an example of the data, the original WebbPSF model, the best-fit model we find with photutils and the residual extended emission for one wavelength close to the [NeĀ III]Ā 15.56 emission, where both point-like source and extended emission are clearly visible (see upper left panel). The lack of PSF structure in the residual map and the good match between the radial distributions within Ā px from the point source location (Channel 3C PSF FWHM Ā px) validate our point-source subtraction procedure.
Fig.Ā 12 shows the ratio between the original and the PSF subtracted emission-line maps, to highlight which regions and emission lines are mainly affected by the contribution of the point source (darker red). The fitted emission lines are shown in order of their wavelength (see Tab.Ā 1). Overall, the point source emission seems to be located in SSC1, as we discuss in Sec.Ā 5.1 (see also Fig.Ā 1).
We applied the same point-source subtraction procedure on the nine band datacubes (A, B, C) of Channel 1, 2 and 3, to retrieve the point-source subtracted spectra for the extended emission continuum and the point source continuum. We used these data to model the continuum with CAFE and estimate the dust attenuation, as explained in Sec.Ā 3.2.
Appendix B Other MIR diagnostic diagrams
In this section we show the [NeĀ III]/[NeĀ II]Ā vs [SĀ IV]/[NeĀ II], color-coded as a function of [OĀ IV]/[NeĀ III]Ā (Fig.Ā 13) and [NeĀ III]/[NeĀ II]Ā versus [OĀ IV]/[NeĀ III], color-coded as a function of [NeĀ V]/[NeĀ II]Ā (Fig.Ā 14), discussed in Sec.Ā 4.4. In the figures, we show the same grids shown in Fig.Ā 8, apart from in Fig.Ā 14 where we also show the R22 disk-plaw models (gray scale) that can reach slightly higher [OĀ IV]/[NeĀ III]. We comment in detail these figures in Sec.Ā 4.4 and Sec.Ā 5.
We highlight that Fig.Ā 13 diagnostic diagram can be affected by dust attenuation since the [SĀ IV]Ā line lies in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, Fig.Ā 3 and Fig.Ā 4). We applied a dust reddening correction (Sec.Ā 3.2) that makes the log([SĀ IV]/[NeĀ II]) line ratio Ā dex higher. Also, this diagram is unable to well-separate the low/intermediate and high-ionization emission, given that [SĀ IV]Ā and [NeĀ III]Ā have similar I.P. and behave similarly. However, it allows us to compare with models also the regions of SBSĀ 0335-052Ā E with no [NeĀ V]Ā or [OĀ IV]Ā emission, that are not shown in Fig.Ā 8 and Fig.Ā 14. The magenta line shown in Fig.Ā 13 reports the empirical relation found by Groves etĀ al. (2008), valid - with some scatter - for very different galaxies, including starbursts, ultra-luminous infrared galaxies, AGN, and also BCDs. Interestingly, SSP and SXPs models seem able to replicate the magenta line only at the highest [SĀ IV]/[NeĀ II], while shocks can only account for the lowest [SĀ IV]/[NeĀ II]Ā and [NeĀ III]/[NeĀ II]. This means that according to the explored set of models, many BCDs could be reproduced only with IMBH models, as shown for SBSĀ 0335-052Ā E.
Fig.Ā 14 diagnostic diagram can be more affected by the wavelength-dependent MIRI/MRS PSF and residual fringing than Fig.Ā 8 and Fig.Ā 13, given that [NeĀ III]Ā lies in Channel 3C and [OĀ IV]Ā in Channel 4C. The effect of the PSF is also probably affecting the ULX [OĀ IV]/[NeĀ III]Ā (magenta diamond), obtained by fitting the ULX spectrum extracted with a variable (PSF-dependent) aperture (see Sec.Ā 3.1). Indeed, the ULX [OĀ IV]/[NeĀ III]Ā looks Ā dex higher than the median value of the spatially resolved data. R22 IMBH low-density models with a low AGN fraction (%) and partially overlap with the data, but are not able to explain the highest [OĀ IV]/[NeĀ III]Ā ratios. Young massive stars may be able to explain the point-source (star symbol) line ratios but struggle to reproduce its [NeĀ V]/[NeĀ II]Ā upper limit as well as its [SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā line ratios shown in Fig.Ā 13.
Appendix C The optical counterparts
Fig.Ā 15 top panel shows the region around the MIRI/MRS point-source (near SSCs1-2; in black) and the UV and [NeĀ V]Ā emitting region (around SSCs4-5; in gray) spectra extracted from the MUSE data (IDĀ 096.B-0690A; PI Hayes), with a 3Ā px (ā) extraction radius at the peak of the HĀ and HeĀ IIĀ 4686 emission, respectively. Consistently with our MIRI/MRS findings, the black spectrum has a larger dust attenuation than the gray one, looking at the different continuum slopes. Interestingly, the HĀ line shows a broad profile in both spectra. This is more evident in the black spectrum, which shows also a broad (fainter) HĀ and [OĀ III]Ā 5007.
Fig.Ā 15 bottom panels show a zoom around the HĀ and [OĀ III]Ā 4959,5007 (left) and HĀ and [NeĀ II]Ā (right) emission lines, with our best-fit shown in red and blue. To reproduce the black and spectra HĀ profiles we added up to 4 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively. To fit the forbidden [OĀ III]Ā line we added up to 3 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
The full modeling and understanding of the complex kinematics highlighted in the MUSE datacube are beyond the scope of this paper. However, the need for (even broad) multiple Gaussian components in spectra located in different parts of the galaxy and in both permitted and forbidden lines indicates that there is no BH broad line region.

### v2

@@
 Abstract
-Nearby blue compact dwarf galaxies (BCDs) share similar properties with objects from the Epoch of Reionization revealed by JWST, in terms of low stellar mass, low metallicity and high specific star-formation rate.
-Thus, they represent ideal local laboratories for detailed multi-wavelength studies to understand their properties and the mechanisms shaping them.
+Nearby blue compact dwarf galaxies (BCDs) are considered analogs to objects from the Epoch of Reionization revealed by JWST, having similar low stellar mass, low metallicity, and high specific star-formation rate.
+Thus, they represent ideal local laboratories detailed for multi-wavelength studies of their properties and mechanisms that shape them.
 We report the first JWST MIRI/MRS observations of the BCD SBSĀ 0335-052Ā E, analyzing MIR emission lines tracing different levels of ionization (e.g., [NeĀ II], [SĀ IV], [NeĀ III], [OĀ IV], [NeĀ V]) of the ionized gas.
 SBSĀ 0335-052Ā E MIR emission is characterized by a bright point source, located in one of the youngest and most embedded stellar clusters (Ā Myr, ), and underlying extended high-ionization emission (i.e., [OĀ IV], [NeĀ V]) from the surroundings of the older and less dusty stellar clusters (Ā Myr, ).
-From the comparison with state-of-the-art models, we can exclude shocks, X-ray binaries, and old stellar populations as the main sources of the high ionization.
-Interestingly, a 4-16% contribution of a Ā Mā intermediate massive black hole (IMBH) is needed to justify the strong [NeĀ V]/[NeĀ II]Ā and would be consistent with optical/UV line ratios from previous studies.
+From the comparison with state-of-the-art models, we can exclude shocks, X-ray binaries, and old stellar populations as the main sources of ionization.
+Interestingly, a 4-8% contribution of a Ā Mā intermediate massive black hole (IMBH) is needed to justify the strong [NeĀ V]/[NeĀ II]Ā and would be consistent with optical/UV line ratios from previous studies.
 However, even IMBH models cannot explain the strongest [OĀ IV]/[NeĀ III].
@@
 In this context, nearby galaxies can provide a precious opportunity for us to understand the characteristics of the interstellar medium (ISM) of these primeval objects and how to distinguish between star-formation and AGN activity within them, given the high signal-to-noise (S/N) of their spectra and the possibility to perform spatially resolved and multi-wavelength studies.
-In particular, nearby blue compact dwarf galaxies (BCDs) are ideal analogs to these high- systems, as they share properties in terms of stellar mass, intense star formation activity and metal poor content, as also suggested by recent JWST studies (e.g., Schaerer etĀ al.
+In particular, nearby blue compact dwarf galaxies (BCDs) are ideal analogs to these high- systems, as they share similar properties in terms of stellar mass, intense star formation activity, and poor metal content, as also suggested by recent JWST studies (e.g., Schaerer etĀ al.
 2022; Brinchmann 2023).
@@
 2024).
-In this context, the mid-infrared (MIR) wavelength range covers a large range of ionization potentials for the same species (e.g., 20-126Ā eV, [NeĀ II]Ā 21.56m, [NeĀ III]Ā 15.56m, [NeĀ V]Ā 14.32,24.3m, [NeĀ VI]Ā ) at lower excitation energies.
+In this context, the mid-infrared (MIR) wavelength range covers a large range of ionization potentials for the same species (e.g., 20-126Ā eV, [NeĀ II]Ā 21.56m, [NeĀ III]Ā 15.56m, [NeĀ V]Ā 14.32,24.3m, [NeĀ VI]Ā 7.65m) at lower excitation energies.
 While these lines are not accessible in the high- universe, they can provide an important constraint on the gas ionization mechanisms and, by also probing other phases (i.e., warm molecular gas with H2 rotational lines; dust with polycyclic aromatic hydrocarbons, PAHs), they can give a broad overview of the ISM conditions.
@@
 2010 and references therein).
-This galaxy is also well-known for a complex kinematics and a large scale HĀ outflow (Herenz etĀ al.
+This galaxy is also well-known for complex kinematics and a large scale HĀ outflow (Herenz etĀ al.
 2023) and high-ionization optical and UV emission lines typically observed in the EoR (e.g., HeĀ IIĀ 4686, CĀ IVĀ , HeĀ IIĀ 1640; Herenz etĀ al.
@@
 2009).
-Nevertheless, MIRI/MRS offer the considerable advantage of spatially resolving the MIR emission (down to ā, Ā pc) and thus revealing the location of very high ionization extended emission (i.e., [NeĀ V]Ā 14.32m) for the first time in BCDs.
+Nevertheless, MIRI/MRS offers the considerable advantage of spatially resolving the MIR emission (down to ā, Ā pc) and thus revealing the location of very high ionization extended emission (i.e., [NeĀ V]Ā 14.32m) for the first time in BCDs.
 In Sec.Ā 2, we present the observations and data reduction.
-In Sec.Ā 3, we show the methods we adopted for our data analysis, including spectral fitting and literature photoionization and shock models that we used to interpret our results, reported in Sec.Ā 4.
+In Sec.Ā 3, we show the methods we adopted for our data analysis, including spectral fitting, and literature photoionization and shock models that we used to interpret our results, reported in Sec.Ā 4.
 In Sec.Ā 5, we discuss our findings as well as possible limitations in current models, while in Sec.Ā 6, we conclude with a summary of our work.
@@
 Dedicated background observations were performed with half the integrations (same group number but half the dithering) to accurately measure and correct for the thermal background.
-Background and science observations were connected via a fixed offset and placed in an uninterruptible sequence, and a PA orientation range was required to obtain simultaneous imaging observations using the F770W, F1130W, and F1800W filters, which will be presented in a forthcoming paper.
-We retrieved MIRI/MRS raw products from Mikulski Archive for Space Telescopes (MAST) and we processed the data with v1.16.0 of the JWST calibration pipeline, following the jupyter notebook provided by the MIRI team.
-As extra steps to the standard version of the MIRI pipeline, we masked additional regions affected by cosmic showers at the end of StageĀ 1 (ShowerMasking Notebook) and we applied residual fringe corrections in StageĀ 2.
+Background and science observations were connected via a fixed offset and placed in an uninterruptible sequence, and a PA orientation range was required to obtain simultaneous imaging observations using the F770W, F1130W, and F1800W filters, which will be presented in a future publication.
+We retrieved MIRI/MRS raw products from Mikulski Archive for Space Telescopes (MAST), and we processed the data with v1.16.0 of the JWST calibration pipeline (Bushouse etĀ al.
+2022), following the jupyter notebook provided by the MIRI team.
+As extra steps to the standard version of the MIRI pipeline, following the recommendations from the JWST pipeline reduction team, we masked additional regions affected by cosmic showers at the end of StageĀ 1 (ShowerMasking Notebook) and we applied residual fringe corrections in StageĀ 2.
 We subtracted the background in StageĀ 3, using the master background created by the pipeline.
-In StageĀ 3 we created the final datacubes per band, to optimize the fringing correction, obtaining 12 datacubes centered to the same coordinates and keeping the native pixel scale size (i.e., 0.13ā, 0.17ā, 0.20ā and 0.35ā, for Channel 1, 2, 3 and 4, respectively).
-Nevertheless, we note that few spaxels located at the peak of the MIR emission are still affected by significant fringing, especially in ChannelsĀ 3 and 4.
+In StageĀ 3, we created the final datacubes per band, to optimize the fringing correction, obtaining 12 datacubes centered to the same coordinates and keeping the native pixel scale size (i.e., 0.13ā, 0.17ā, 0.20ā and 0.35ā, for Channel 1, 2, 3 and 4, respectively).
+Nevertheless, we note that the spaxels surrounding the bright point source that we further discuss below, within a circular region with a radius equal to the MIRI/MRS PSF FWHM, are still significantly affected by fringing, especially in ChannelsĀ 3 and 4.
 The FOV progressively increases from 3.2ā3.7ā in Channel 1, to 6.6ā7.7ā in Channel 4, as well as the MIRI/MRS PSF and spatial resolution, from in Channel 1 to in Channel 4.
-In Fig.Ā 1 we show the position of MIRI/MRS FOV superimposed on the HST ACS F150LP UV emission (left panel) and the 12-17Ā m slice from MIRI/MRS Channel 3 (right panel).
-The UV emission (map and grey contours) clearly shows the six young, massive and compact āsuper star clustersā (SSC1, SSC2, SSC3, SSC4, SSC5, SSC6) revealed in this galaxy (Thuan etĀ al.
+In Fig.Ā 1, we show the position of MIRI/MRS FOV superimposed on the HST ACS F150LP UV emission (left panel) and the 12-17Ā m slice from MIRI/MRS Channel 3 (right panel).
+The UV emission (map on the left panel, gray contours on the right panel) clearly shows the six young, massive and compact āsuper star clustersā (SSC1, SSC2, SSC3, SSC4, SSC5, SSC6) revealed in this galaxy (Thuan etĀ al.
 1997).
@@
 2001).
-The HeĀ IIĀ 4686 high-ionization emission revealed with MUSE (white dotted contours; seeing ā) follows the UV emission and peaks around SSCs4,5.
+The HeĀ IIĀ 4686 high-ionization emission revealed with MUSE (white dotted contour on the right panel; seeing ā) follows the UV emission and peaks around SSCs4,5.
 The HST H black dashed contours highlight the three additional actively young star-forming regions, S3-Pa, S7 and S8, discovered by Thompson etĀ al.
@@
 3.1 Emission-line fitting
-In each spectrum we fitted several MIR emission lines, performing the line fitting locally around each line in a km/s window centered at the observed wavelength of the line of interest.
+In each spectrum, we fitted several MIR emission lines, performing the line fitting locally around each line in a km/s window centered at the observed wavelength of the line of interest.
 In particular, we used a combination of a Gaussian profile and a one-degree polynomial to account for emission-line and surrounding continuum emission in each spaxel of the point-source subtracted band datacube slices, as well as in the 1-D spectrum of the bright point-source and the spectrum extracted at the ULX position (shown in Fig.Ā 1).
-The 1-D spectrum of the bright point source is obtained by the best-fit model of the PSF photometry analysis described in App.Ā A, while the ULX spectrum is extracted from a region with variable diameter depending on the channel, matching the FWHM of the PSF (i.e., Ā 0.4ā, 0.5ā, 0.6ā, 0.9ā (Law etĀ al.
+The 1-D spectrum of the bright point source is obtained by the best-fit model of the PSF photometry analysis described in App.Ā A, while the ULX spectrum is extracted from a region with variable diameter depending on the channel, matching the FWHM of the PSF (i.e., Ā 0.4ā, 0.5ā, 0.6ā, 0.9ā; Law etĀ al.
 2023).
-Table 1 compiles the list of emission lines analyzed, together with their rest-frame wavelengths and ionization potentials (I.P.).
+Tab.Ā 1 compiles the list of emission lines analyzed, together with their rest-frame wavelengths and ionization potentials (I.P.).
 | Line | I.P. | |
@@
 | [OĀ IV] | 25.89 | 54.95 |
-Fig.Ā 2, 3 and 4 show the point source spectrum (SSC1), the spectrum extracted from the spaxels covering the ULX position, and a spectrum from the north-west region covering the clusters SSCs3,4,5,6 and S7 and the peak of UV emission (see Fig.Ā 1), respectively, as well as the best-fit (in red) for the MIR emission lines.
+Fig.Ā 2, 3 and 4 show the point source spectrum (SSC1), the spectrum extracted from the spaxels covering the ULX position, and a spectrum from the north-west region covering the older clusters SSCs3,4,5,6 and S7, the peak of UV emission and the ULX (see Fig.Ā 1), respectively, as well as the best-fit (in red) for the MIR emission lines.
 In the following, we consider line detections as those with fluxes with signal-to-noise (S/N) higher than 3 and upper limits for measurements with S/N2.
 In particular, we note that emission lines typically observed in AGN (e.g., Dasyra etĀ al.
-2024), such as [NeĀ VI]Ā Ā 7.65 (tentatively fitted only in the bright point source spectrum with ), [MgĀ V]Ā Ā 5.61 (I.P.Ā Ā eV) and [MgĀ VII]Ā Ā 5.5 (I.P.Ā Ā eV) are not detected.
+2024), such as [NeĀ VI]Ā Ā 7.65 (tentatively fitted only in the bright point source spectrum but with ), [MgĀ V]Ā Ā 5.61 (I.P.Ā Ā eV) and [MgĀ VII]Ā Ā 5.5 (I.P.Ā Ā eV) are not detected.
 3.2 Dust attenuation
@@
 (2022) and the Gordon etĀ al.
-(2023) dust extinction relationship.
-When exploring the spectrum of the point source SSC1, we noticed a wide absorption in the 7.5-12.5m range, corresponding to the silicate feature at 9.7Ā m (highlighted in grey in Fig.Ā 2, 3 and 4).
+(2023) dust extinction relationship (see also Gordon etĀ al.
+2009; 2021; Decleir etĀ al.
+2022; Fitzpatrick etĀ al.
+2019).
+When exploring the spectrum of the point source SSC1, we noticed a wide absorption in the 7.5-12.5Ā m range, corresponding to the silicate feature at 9.7Ā m (highlighted in gray in Fig.Ā 2, 3 and 4).
 This feature (revealed also in SBSĀ 0335-052Ā E Spitzer spectrum; Houck etĀ al.
@@
 Thus, an intrinsic dust attenuation correction has to be applied to the line fluxes as well.
-Before using the continuum to determine the level of dust attenuation, there are few extra corrections we needed to implement to the datacubes.
+Before using the continuum to determine the level of dust attenuation, there are a few extra corrections we needed to implement to the datacubes.
 First, we limited the impact of fringing on the continuum spectra by applying an additional 1D fringing correction using the JWST pipeline tools.
@@
 With the continuum corrected and optimized for the analysis, we used the Continuum and Feature Extraction software (CAFE, Marshall etĀ al.
-2007; Diaz-Santos et al., in prep.) to model the continuum of both the point-source spectrum and the point-source subtracted (extended) continuum in the range Ā m from Channel 1, 2 and 3 band datacubes (see App.Ā A).
+2007; Diaz-Santos etĀ al.
+2025) to model the continuum of both the point-source spectrum and the point-source subtracted (extended) continuum in the range Ā m from Channel 1, 2 and 3 band datacubes (see App.Ā A).
 CAFE uses a combination of blackbody emission at different temperatures together with a custom attenuation law (compatible with Gordon etĀ al.
-2023) to describe the overall shape of the continuum at near-to-mid infrared wavelengths.
+2023 law; see also Gordon etĀ al.
+2009; 2021; Decleir etĀ al.
+2022; Fitzpatrick etĀ al.
+2019) to describe the overall shape of the continuum at near-to-mid infrared wavelengths.
 For the case of SBS 0335-052Ā E only two dust components were considered: a cool component, whose temperature falls in the range 40-100 K, and a warm component, with temperature in the range 100-500 K.
@@
 We used the derived values and the Gordon etĀ al.
-(2023) dust extinction relationship to correct for intrinsic dust attenuation.
-Although the extinction curve at the MIR wavelengths is almost flat, lines located in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, 3 and 4) can suffer from large attenuation.
+(2023) dust extinction relationship (see also Gordon etĀ al.
+2009; 2021; Decleir etĀ al.
+2022; Fitzpatrick etĀ al.
+2019) to correct for intrinsic dust attenuation.
+Although the extinction curve at the MIR wavelengths is almost flat, lines located in the deepest part of the silicate feature (highlighted in gray in Fig.Ā 2, 3 and 4) can suffer from large attenuation.
 In the case of SBSĀ 0335-052Ā E, the dust-corrected fluxes of almost every emission line are increased by a factor of () in the point-like source (extended) spectrum, except for the [ArĀ III]Ā and [SĀ IV]Ā lines ā which fall within the silicate feature ā whose correction factor is ().
@@
 ā¢
-SSP models with and without the ionizing output from X-ray binaries (XRBs) from Garofali etĀ al. (2024; G24 hereafter);
+SSP models with and without the ionizing output from X-ray binaries (XRBs) from Garofali etĀ al. (2024; G24 hereafter)222https://github.com/kgarofali/sxp-cloudy;
 ā¢
-Intermediate-massive black hole (IMBH) models from Richardson etĀ al. (2022; R22 hereafter);
+Intermediate-massive black hole (IMBH) models from Richardson etĀ al. (2022; R22 hereafter)333https://github.com/crichardson17/richardson_2022;
 ā¢
-Shock models from Flury etĀ al. (2024; F24 hereafter) and the Mexican Million Models Shock Database222http://3mdb.astro.unam.mx:3686/, described in Alarie & Morisset (2019; AM19 hereafter).
+Shock models from Flury etĀ al.
+(2024; F24 hereafter)444https://zenodo.org/records/14140949 and the Mexican Million Models Shock Database555http://3mdb.astro.unam.mx:3686/, described in Alarie & Morisset (2019; AM19 hereafter).
 For all these models, we considered only the two lowest metallicity grids available (i.e., 5% and 10% Zā models for G24, R22, F24, and 2% and 10% Zā for MP23), in accordance with SBSā0335-052Ā Eās metallicity range (Papaderos etĀ al.
@@
 For MP23, R22 and shock models, we could also test the effect of different densities (Ā Ā cm-3 in MP23, R22; Ā Ā cm-3 in shock models), finding small differences that only slightly impact the results shown in Sec.Ā 4, as we discuss in Sec.Ā 5.
-Here below we give a summary of the characteristics of all these models.
+In the following, we give a summary of the characteristics of all these models.
 MP23 SSP models combine CLOUDY (Ferland etĀ al.
 2017) with the revised version of the Bruzual & Charlot (2003) stellar population synthesis models introduced in Plat etĀ al.
-(2019; C&B models), spanning across several parameters (see their TableĀ 4), including a Kroupa (2001) and top-heavy (x030) initial mass functions (IMFs) with a mass range up to 100 and 300 Mā.
+(2019; C&B models), spanning across several parameters (see their Tab.Ā 4), including a Kroupa (2001) and top-heavy (x030) initial mass functions (IMFs) with a mass range up to 100 and 300 Mā.
 C&B models accurately cover the evolution of O and B stars, including the Wolf-Rayet (WR) phase, whose contributions last a few Myr.
@@
 In particular, in this paper we show the 20Ā Myr instantaneous burst of star formation grids (in SSPs [OĀ III]Ā 5007/HĀ peaks at 20Ā Myr; Xiao etĀ al.
-2018), with coincident mixing and closed geometry.
+2018), with coincident mixing and different geometry.
 The line ratios do not vary significantly in the non-coincident outputs, but can be slightly affected by the choice of geometry, as we comment in Sec.Ā 4.
@@
 In this work, we consider only shocks+precursor models, since they also consider the gas entering the shock front.
-The grids we show span values of shock velocity (Ā km/s) and the magnetic field parameter (Ā Gauss), which are the main parameters regulating the shock ionization spectrum and the effective ionization parameter.
+The shown grids span values of shock velocity (Ā km/s) and the magnetic field parameter (Ā Gauss), which are the main parameters regulating the shock ionization spectrum and the effective ionization parameter.
 We find that AM19 and F24 shock models behave similarly in the range of parameters considered, thus in Sec.Ā 4 we show only F24 grids.
@@
 The FOV varies from 3.2ā3.7ā, to 4.0ā4.8ā, 5.2ā6.2ā and 6.6ā7.7ā for Channel 1 (Pf), 2 ([ArĀ III], [SĀ IV]), 3 (Hu, [NeĀ II], [NeĀ V], [NeĀ III]) and 4 ([SĀ III], [OĀ IV]), respectively.
-In the [SĀ III]Ā map panel we indicate the positions of SBSĀ 0335-052Ā E several star-forming regions as we did in Fig.Ā 1.
+In the [SĀ III]Ā map panel, we indicate the positions of several star-forming regions in SBSĀ 0335-052Ā E as we did in Fig.Ā 1.
 PfĀ and HuĀ maps follow the HĀ distribution, which peaks in young star-forming regions (Ā Myr), such as the position of SSC1 (point source) and SSC2 as well as S7 and S3-Pa, slightly shifted with respect to the other SSCs and very bright in Pa (Thompson etĀ al.
@@
 We notice that the peak of [NeĀ III]Ā and [SĀ IV]Ā emission is close to the ULX position revealed by Prestwich etĀ al.
-(2013; see also ) with Chandra (magenta circle).
+(2013; their TableĀ 3; see also ) with Chandra (magenta circle).
 However, there is no clear correspondence between the ULX X-ray emission and high ionization lines, since the [NeĀ V], [OĀ IV]Ā and UV emission peaks are shifted towards S7.
 We highlight that the ULX extracted spectrum is also characterized by [FeĀ II]Ā 5.34m emission, that is not shown in Fig.Ā 5 since it does not show clear extended emission (see Fig.Ā 12).
-The region where we subtracted the point-source emission, located at the position of SSC1, shows either a reduced or tentative emission (grey regions with ) in the line maps.
+The region where we subtracted the point-source emission, located at the position of SSC1, shows either a reduced or tentative emission (gray regions with ) in the line maps.
 This can be clearly seen in the HuĀ and [NeĀ V]Ā maps.
@@
 As shown in Fig.Ā 2 the point source SSC1 shows almost all the emission lines reported in Fig.Ā 5, apart from [NeĀ V]Ā and [ArĀ II].
+Finally, we extracted a spectrum centered at the point source position in the residual cube (i.e., after point-source subtraction), summing all the spaxels within one PSF FWHM, confirming that the absence of [NeĀ V]Ā is not due to residual fringing in ChannelĀ 3 (see SecĀ 2).
 The point source spectrum also shows clear [FeĀ II]Ā emission (see also Fig.Ā 12).
 4.2 Emission-line ratio maps
-Fig.Ā 6 shows [NeĀ III]/[NeĀ II], [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ II]Ā line ratio extended emission maps with a color-coding chosen to enhance the peak regions to aid with the interpretation of the diagnostic diagrams.
+Fig.Ā 6 shows [NeĀ III]/[NeĀ II], [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā line ratio extended emission maps with a color-coding chosen to enhance the peak regions to aid with the interpretation of the diagnostic diagrams.
 [NeĀ III]/[NeĀ II]Ā is known to be a good tracer of the ionization parameter log(), while [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā are found to be sensitive also to the hardness of the ionizing spectrum (Richardson etĀ al.
@@
 Finally, S8 does not have either [NeĀ V]Ā or [OĀ IV]Ā emission but shows intermediate values of [NeĀ III]/[NeĀ II].
+Tab.Ā 2 reports the line ratios for the different regions discussed above.
 4.3 Ionized gas kinematics maps
@@
 The velocity maps are centred on the systemic velocity of this galaxy according to previous optical studies (Ā km/s; Moiseev etĀ al.
-2010)333 We centred the [NeĀ III]Ā velocity map -24Ā km/s below the systemic velocity Ā km/s found by Moiseev etĀ al.
+2010)666 We centred the [NeĀ III]Ā velocity map -24Ā km/s below the systemic velocity Ā km/s found by Moiseev etĀ al.
 (2010).
@@
 2017; 2023).
-In particular, [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā maps trace progressively higher ionization gas (Tab.Ā 1) and show a consistent velocity pattern, with a blue-shifted emission around the position of S3-Pa and in the north-west portion of the MIRI/MRS FOV (on the right of S7 and at the top of S8), and red-shifted emission from north to south and in the east region.
+In particular, [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā maps trace progressively higher ionization gas (Tab.Ā 1) and show a consistent velocity pattern, with a blueshifted emission around the position of S3-Pa and in the north-west portion of the MIRI/MRS FOV (on the right of S7 and at the top of S8), and redshifted emission from north to south and in the east region.
 Despite the kinematic complexity shown in the velocity maps, the emission lines are well-fitted with a single Gaussian, broadening in the north-west region of the MIRI/MRS FOV.
@@
 We notice that the region with enhanced in the [NeĀ V]Ā map is located close to the peak of the [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ II]Ā line ratios (Fig.Ā 6) tracing the hardening of the radiation field.
-The regions where [NeĀ III]Ā and [SĀ IV]Ā peaks probably have no [NeĀ V]Ā emission due to their lower ionization (they also partly show [NeĀ II]Ā emission; Fig.Ā 5) and are characterized by enhanced red-shifted velocities, possibly showing a flow of gas moving away from the [NeĀ V]Ā high ionization region.
+The regions where [NeĀ III]Ā and [SĀ IV]Ā peaks probably have no [NeĀ V]Ā emission due to their lower ionization (they also partly show [NeĀ II]Ā emission; Fig.Ā 5) and are characterized by enhanced redshifted velocities, possibly showing a flow of gas moving away from the [NeĀ V]Ā high ionization region.
 This picture is consistent with what found by Izotov etĀ al.
@@
 [SĀ IV]/[NeĀ II]Ā instead can be affected by dust reddening, while [OĀ IV]/[NeĀ III]Ā is the ratio most affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing, given that [NeĀ III]Ā is located in ChannelĀ 3c and [OĀ IV]Ā in ChannelĀ 4c.
+This is also likely affecting the [OĀ IV]/[NeĀ III]Ā value estimated for the ULX extracted spectrum, that is Ā dex higher than the median value of the spatially resolved spaxels (see Fig.Ā 14, Tab.Ā 2).
 For these reasons, we mainly focus this section on [NeĀ V]/[NeĀ II]Ā versus [NeĀ III]/[NeĀ II], moving the other two diagnostic diagrams in App.Ā B.
@@
 These two grids are the only ones (partially) able to cover the x- and y- ranges shown.
-However, we discarded the 100Ā Myr grid because old stars struggle to explain the observed number of ionizing photons observed in this system as we discuss in Sec.Ā 5.2.1.
+However, we discarded the 100Ā Myr grid because old stars struggle to explain the observed number of ionizing photons observed in this system, as we discuss in Sec.Ā 5.2.1.
 Also, we notice that the choice of the IMF does not affect the line ratios, but the choice of Ā Mā strongly enhances [NeĀ V]/[NeĀ II].
@@
 However, we highlight that these IMBH models still struggle to reproduce the highest [OĀ IV]/[NeĀ III]Ā line ratios (Fig.Ā 14).
-Literature BCDs [SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā observed with Spitzer follow the Groves etĀ al.
-2008 empirical relation shown in Fig.Ā 13 (magenta line).
-This indicates that low-metallicity SSP, SXP and shock grids generally struggle to reproduce BCDsā lower ionization MIR lines than [NeĀ V].
+[SĀ IV]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā ratios observed with Spitzer in BCDs are consistent with the empirical relation from Groves etĀ al.
+2008, shown as the magenta line in Fig.Ā 13.
+Since low-metallicity SSP, SXP, and shock models fail to reproduce this relation, they also struggle to match the lower-ionization MIR lines (below [NeĀ V]) typically observed in BCDs.
 Given the critical densities of the MIR lines, another property that can affect the parameter space covered by these line ratios is the gas density.
@@
 | SSC2 | 3.0 | 3.7 | SSC1-2 uniquely has dust and H2, and thermal radio emission (J09) | ||
-SSC3(+S3-Pa) | 7.0 | 7.1 | 7.31 | WR Blue bump (P06; K18) | |
+| SSC3 | 7.0 | 7.1 | 7.31 | WR Blue bump (P06; K18) | |
 | SSC4 | 11.0 | 1.1 | 7.27 | SSCs4-5 characterized by ([NeĀ V]) peak and tentative radio emission (J09) | |
@@
 | SSC6 | 11.0 | 2.6 | ā | Point source in NICMOS 1.6Ā m (T09); ([SĀ IV]) and ([NeĀ III]) peak | |
-S3-Pa | (?) | ā | 7.38 | Bright H lines (T09) | |
-S7 | (?) | ā | 7.10 | Bright H lines (T09); [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā peak | |
-S8 | ā | ā | 7.21 | ||
+| S3-Pa | Ā (T09) | 2.3Ā (T09) | 7.38 | Bright H lines (T09) | |
+| S7 | Ā (T09) | 9.4Ā (T09) | 7.10 | Bright H lines (T09); [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā peak | |
+| S8 | Ā (T09) | 2.2Ā (T09) | 7.21 | ||
 | ULX | ā | ā | ā | Revealed by P13 (see also K18) | |
 | Region | [NeĀ III]/[NeĀ II] | [NeĀ V]/[NeĀ II] | [OĀ IV]/[NeĀ III] | [SĀ IV]/[NeĀ II] | [NeĀ V]Ā 14.3m |
-SSC1 | |||||
+| SSC1ā | |||||
 | SSC2 | |||||
@@
 | S8 | ā | ā | |||
-ULX | ā |
+| ULXā | ā | ||||
 5.1 The bright point source, SSC1
 As introduced in Sec.Ā 3, the point source is the source SSC1 (Fig.Ā 1).
-SSC1 appears as a point source in the MIRI/MRS data possibly because of its dusty nature, which makes it particularly bright in the MIR.
+SSC1 appears as a point source in the MIRI/MRS data, possibly because of its dusty nature, which makes it particularly bright in the MIR.
 As described in Sec.Ā 3.2, when analyzing the MIR continuum, we find a very high attenuation for the point source spectrum ().
@@
 In particular, it is important to consider that [OĀ IV]Ā has a similar I.P.
-to HeĀ IIĀ (Ā ryd), while [NeĀ V]Ā requires the presence of even harder radiation (Ā ryd), corresponding to the extreme ultraviolet and soft X-ray range.
+to HeĀ II, while [NeĀ V]Ā requires the presence of even harder radiation, corresponding to the extreme ultraviolet and soft X-ray range (see Tab.Ā 1).
 5.2.1 Pure Bursts of Star Formation
@@
 As mentioned previously in Sec.Ā 5.1, Kehrig etĀ al.
-(2018) was only able to derive SBSĀ 0335-052Ā Eās HeĀ II-ionizing budget via either single, rotating metal-free massive stars (Ā MĀ Ā Mā stars, see their Sec.Ā 5.3.1) or a binary population with 0.05%Ā Zā metallicity and a top-heavy IMF (modeled with BPASS; see their Sec.Ā 5.3.2).
+(2018) was only able to derive SBSĀ 0335-052Ā Eās HeĀ II-ionizing budget via either single, rotating metal-free massive stars (see their Sec.Ā 5.3.1) or a binary population with 0.05%Ā Zā metallicity and a top-heavy IMF (modeled with BPASS; see their Sec.Ā 5.3.2).
 However, a 0-0.05%Ā Zā metallicity is unphysically low for SBSĀ 0335-052Ā E (see also Wofford etĀ al.
@@
 2006; Kehrig etĀ al.
-2018), the large number needed to justify the HeĀ II-ionization budget () would overcome by a factor of the SSCs total mass, depending on the IMF choice (Kehrig etĀ al.
+2018), the large number needed to justify the HeĀ II-ionization budget would overcome by a factor of the SSCs total mass, depending on the IMF choice (Kehrig etĀ al.
 2018).
@@
 Wofford etĀ al.
-(2021; see their SecĀ 3.1) proposed the presence of very massive stars (VMSs) - stars with masses above 100Ā Mā (Vink 2012) - from the blue-shifted [OĀ V]Ā 1371 detection in HST COS data, covering a 2.5ā circular region centered on the peak of the UV emission (see Fig.Ā 1).
+(2021; see their SecĀ 3.1) proposed the presence of very massive stars (VMSs) - stars with masses above 100Ā Mā (Vink 2012) - from the blueshifted [OĀ V]Ā 1371 detection in HST COS data, covering a 2.5ā circular region centered on the peak of the UV emission (see Fig.Ā 1).
 Thus, one might suspect that VMSs might represent the definite answer, hardening the radiation field and enabling highly ionized extended emission.
@@
 Unfortunately, the shell structure clearly visible in the HeĀ IIĀ morphology is not covered in our MIRI/MRS observations.
-This region is slightly north-west to the area where we see Ā km/s red-shifted velocities and intrinsic velocity dispersion enhancement in [SĀ IV]Ā and [NeĀ III]Ā (Fig.Ā 7).
+This region is slightly north-west to the area where we see Ā km/s redshifted velocities and intrinsic velocity dispersion enhancement in [SĀ IV]Ā and [NeĀ III]Ā (Fig.Ā 7).
 Having said that, as shown in Figs.Ā 8, 13 and 14, low-metallicity F24 (and also AM19, not shown) shock models do not overlap with our data.
-The grids could match the range of observed [OĀ IV]/[NeĀ III], but cannot explain the corresponding [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios.
+The grids could match the range of observed [OĀ IV]/[NeĀ III]Ā (and other high-ionization line-ratios, such as [NeĀ V]/[NeĀ III]Ā and [OĀ IV]/Pf), but cannot explain the corresponding [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios.
+Overall, shock models seem to overpredict low-ionization lines such as [NeĀ II].
 Another hint towards shocks not being the primary source of ionization of the extended [OĀ IV]Ā and [NeĀ V]Ā emission is the complete lack of co-spatial [FeĀ II]Ā emission (see Sec.Ā 5.1).
@@
 Hatano etĀ al.
-(2023) proposed also the presence of broad HĀ emission possibly due to the BH broad line region, from which they estimate a Ā Mā upper limit.
+(2023) also proposed the presence of broad HĀ emission possibly due to the BH broad line region, from which they estimate a Ā Mā upper limit.
 However, this value is much larger than the mass of the different components of this galaxy (see Tab.Ā 2), raising doubts that this emission is tracing the BH broad line region (see also Sec.Ā 5.3), unless the system is not virialized (e.g., Bertemes etĀ al.
@@
 As such, the existence of [NeĀ V]Ā 3426 and the observed NIR variability seem the two most feasible clues from the literature to suggest the presence of an IMBH in SBSĀ 0335-052Ā E.
-In consideration of these suggestions of an IMBH, looking at Fig.Ā 8 and Fig.Ā 13, the models that uniquely cover the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios that we measure in SBSĀ 0335-052Ā E are the Ā Mā IMBH models from R22 with an AGN fraction 4-8-16% and log() between and (the closed geometry model can better reach the highest [NeĀ III]/[NeĀ II]).
+In consideration of these suggestions of an IMBH, looking at Fig.Ā 8 and Fig.Ā 13, the models that uniquely cover the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios that we measure in SBSĀ 0335-052Ā E are the Ā Mā IMBH models from R22 with an AGN fraction 4-8% and log() between and (the closed geometry model can better reach the highest [NeĀ III]/[NeĀ II]).
 According to the R22 models, we can exclude the possibility of a BH with Ā Mā because this BH mass would imply too low log([NeĀ III]/[NeĀ II]) () and log([NeĀ V]/[NeĀ II]) () with respect to the observed data, regardless of the other parameter selection (i.e., AGN fraction, geometry, gas density, stellar age).
@@
 2023).
-In particular, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā intrinsic velocity dispersion maps revealed an interesting āV shapeā in the north-west region, starting at the position of SSCs 4-5, high-resolution optical IFU data revealed doubled-peak HĀ and HeĀ IIĀ profiles, suggesting the presence of an expanding shell (Izotov etĀ al.
+In particular, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā velocity dispersion (corrected for instrumental broadening) maps revealed an interesting āV shapeā in the north-west region, starting at the position of SSCs 4-5, high-resolution optical IFU data revealed doubled-peak HĀ and HeĀ IIĀ profiles, suggesting the presence of an expanding shell (Izotov etĀ al.
 2006).
@@
 However, we do stress that even the R22 models do not perfectly overlap with our data, since they are not able to reproduce the combination of highest [OĀ IV]/[NeĀ III]Ā and [NeĀ III]/[NeĀ II]Ā values (see Fig.Ā 14).
-Many factors could be at play here with regards to the assumptions and parameters held within the models - each of which are difficult to constrain with observations alone due to the degeneracies between them.
+Many factors could be at play here with regard to the assumptions and parameters held within the models - each of which are difficult to constrain with emission-line ratio diagrams alone due to the degeneracies between them.
 For instance, R22 predictions can vary as a function of the considered SEDs (R22 considered two extreme values, qso or disk-plaw) and assumptions on accretion rate and spin (incorporated as fixed parameters in the qso SED), as well as the choice of geometry (open vs closed).
@@
 C).
+This implies that it may have an extended nature, not compatible with a BH broad line region.
 In particular, both the H and [O III] 5007 emission lines require up to 3-4 Gaussian components to be well fitted, with also [O III] 5007 showing a fainter broad (up to FWHM  km/s) component.
@@
 Exploiting the use of high- analogs, Mingozzi etĀ al.
-(2024) used optical and UV spectra from the Cosmic Origins Spectrograph (COS) Legacy Archive Spectroscopic SurveY (CLASSY; Berg etĀ al.
+(2024) used narrow lines in optical and UV spectra from the Cosmic Origins Spectrograph (COS) Legacy Archive Spectroscopic SurveY (CLASSY; Berg etĀ al.
 2022; James etĀ al.
@@
 Specifically, Mingozzi etĀ al.
-(2024) proposed the CĀ III]Ā 1907,9/HeĀ IIĀ 1640 versus OĀ III]Ā 1666/HeĀ IIĀ 1640 as the most reliable UV diagnostic diagram (their Fig.Ā 6), being capable of separating the SF, AGN and shocks at low-metallicity.
-We re-propose this diagnostic diagram in Fig.Ā 9, showing the CLASSY sample compared with R22 models with the same parameters presented in FigureĀ 8 and discussed in SectionĀ 5.2.4.
-The fact that SBSĀ 0335-052Ā E (light-pink square) lies very close to the line separators was interpreted by Mingozzi etĀ al.
+(2024) proposed the CĀ III]Ā 1907,9/HeĀ IIĀ 1640 versus OĀ III]Ā 1666/HeĀ IIĀ 1640 as the most reliable UV diagnostic diagram (their Fig.Ā 6; see also Feltre etĀ al.
+2016), being capable of separating the SF, AGN and shocks at low-metallicity.
+We revisit this diagnostic diagram in Fig.Ā 9, showing the CLASSY sample compared with R22 models with the same parameters presented in FigureĀ 8 and discussed in SectionĀ 5.2.4.
+The fact that SBSĀ 0335-052Ā E (light-green square) lies very close to the line separators was interpreted by Mingozzi etĀ al.
 (2024) as being due a combination of its low-metallicity environment and hard stellar radiation field.
@@
 Since the UV emission line fluxes are from spectra integrated over the COS aperture (2.5ā; see Fig.Ā 1) covering the [NeĀ V]Ā 14.3 emitting region, the fact that the UV line ratios can be reproduced by the same set of models that explain part of the spatially resolved MIR line ratios indeed strengthens the validity of the R22 models in describing this object.
-This finding also highlights a previously unknown degeneracy of this diagnostic diagram in distinguishing between pure star-forming objects and objects with an AGN fraction Ā % (when considering R22 models).
+This finding also highlights a previously unknown degeneracy of this diagnostic diagram in distinguishing between pure star-forming objects and objects with an AGN fraction Ā % (when considering R22 models with Ā Mā).
 The MIR wavelength range is not observable in high- targets, but JWST does provide coverage of [NeĀ V]Ā in the optical wavelength range (e.g., Cleri etĀ al.
@@
 Consistently with the MIR (Fig.Ā 8) and UV line ratios (Fig.Ā 9), SBSĀ 0335-052Ā Eās optical line ratios are also consistent with the R22 Ā Mā IMBH models with 8% AGN fraction.
-Interestingly, SBSĀ 0335-052Ā Eās [NeĀ V]/[NeĀ III]Ā flux ratio (pink square) is below the median values found for the 13 currently known nearby BCD [NeĀ V]Ā emitters (Izotov etĀ al.
+Interestingly, SBSĀ 0335-052Ā Eās [NeĀ V]/[NeĀ III]Ā flux ratio (light-green square) is below the median values found for the 13 currently known nearby BCD [NeĀ V]Ā emitters (Izotov etĀ al.
 2004; 2012; 2021; Thuan & Izotov 2005) and is of the value found for the galaxy studied by Chisholm etĀ al.
@@
 We modelled the PSF for each MIRI/MRS band using vĀ 1.4.0 WebbPSF (Perrin etĀ al.
-2014) and the function calc_datacube444https://webbpsf.readthedocs.io/en/latest/jwst_ifu_datacubes.html, taking into account the same native pixel scale and spectral binning of the observed data.
+2014) and the function calc_datacube777https://webbpsf.readthedocs.io/en/latest/jwst_ifu_datacubes.html, taking into account the same native pixel scale and spectral binning of the observed data.
 We created the pipeline Stage 3 products by band, selecting the option ifualign, to have them aligned with the WebbPSF models.
@@
 We comment in detail these figures in Sec.Ā 4.4 and Sec.Ā 5.
-We highlight that Fig.Ā 13 diagnostic diagram can be affected by dust attenuation since the [SĀ IV]Ā line lies in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, Fig.Ā 3 and Fig.Ā 4).
+We highlight that Fig.Ā 13 diagnostic diagram can be affected by dust attenuation since the [SĀ IV]Ā line lies in the deepest part of the silicate feature (highlighted in gray in Fig.Ā 2, Fig.Ā 3 and Fig.Ā 4).
 We applied a dust reddening correction (Sec.Ā 3.2) that makes the log([SĀ IV]/[NeĀ II]) line ratio Ā dex higher.
@@
 This is more evident in the black spectrum, which shows also a broad (fainter) HĀ and [OĀ III]Ā 5007.
-Fig.Ā 15 bottom panels show a zoom around the HĀ and [OĀ III]Ā 4959,5007 (left) and HĀ and [NeĀ II]Ā (right) emission lines, with our best-fit shown in red and blue.
-To reproduce the black and spectra HĀ profiles we added up to 4 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
-To fit the forbidden [OĀ III]Ā line we added up to 3 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
+Fig.Ā 15 bottom panels show a zoom around the HĀ and [OĀ III]Ā 4959,5007 (left) and HĀ and [NĀ II]Ā (right) emission lines, with our best-fit shown in red and blue.
+To reproduce the black and gray spectra HĀ profiles we added up to 4 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
+To fit the forbidden [OĀ III]Ā 5007 line in the black and gray spectra we added up to 3 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
 The full modeling and understanding of the complex kinematics highlighted in the MUSE datacube are beyond the scope of this paper.
-However, the need for (even broad) multiple Gaussian components in spectra located in different parts of the galaxy and in both permitted and forbidden lines indicates that there is no BH broad line region.
+However, the need for (even broad) multiple Gaussian components in spectra located in different parts of the galaxy and in both permitted and forbidden lines suggests that there is no BH broad line region.
+Similar broad components (Ā km/s) have been reported in other metal-poor galaxies (e.g., the brightest star-forming region in CGCG 007-025, del Valle-Espinosa etĀ al.
+2023), whose extension also prevents its association with a BH broad line region.